### v1

# Evaluating and Improving the Robustness of LiDAR-based Localization and Mapping

Abstract
LiDAR is one of the most commonly adopted sensors for simultaneous localization and mapping (SLAM) and map-based global localization. SLAM and map-based localization are crucial for the independent operation of autonomous systems, especially when external signals such as GNSS are unavailable or unreliable. While state-of-the-art (SOTA) LiDAR SLAM systems could achieve 0.5% (i.e., 0.5m per 100m) of errors and map-based localization could achieve centimeter-level global localization, it is still unclear how robust they are under various common LiDAR data corruptions.
In this work, we extensively evaluated five SOTA LiDAR-based localization systems under 18 common scene-level LiDAR point cloud data (PCD) corruptions. We found that the robustness of LiDAR-based localization varies significantly depending on the category. For SLAM, hand-crafted methods are in general robust against most types of corruption, while being extremely vulnerable (up to +80% errors) to a specific corruption. Learning-based methods are vulnerable to most types of corruptions. For map-based global localization, we found that the SOTA is resistant to all applied corruptions. Finally, we found that simple Bilateral Filter denoising effectively eliminates noise-based corruption but is not helpful in density-based corruption. Re-training is more effective in defending learning-based SLAM against all types of corruption.
I Introduction
Robotics and automation systems have been gaining significant prevalence in this decade. Of those, robots, drones, and autonomous vehicles have been making the news [1] due to their widespread adoption and related safety concerns [2].
Localization is fundamental to these systems since it keeps track of the position of a robotic system within an environment. Outdoors, and Global Navigation Satellite Systems (GNSS) are widely used for localization. In contrast, in urban, tunnel, and indoor settings where GNSS is hindered by weak satellite signals [3], simultaneous localization and mapping (SLAM) systems are favored since they operate independently of external signals and reliably navigate both indoors and outdoors. For these reasons, SLAM is widely used for locating and mapping-related tasks in robotics and autonomous driving. SLAM methods utilize either a camera, LiDAR, or a combination of both, with the optional integration of an Inertial Measurement Unit (IMU) [4]. Compared to vision SLAM, LiDAR SLAM has higher mapping accuracy, and better stability, is less susceptible to lighting conditions, and does not observe scale drift. These advantages lead to its popularity in large-scale use cases and facilitate more research and application in this direction.
Despite the recent successes of LiDAR SLAM, many challenges remain. Laser signals are robust under clear conditions, however, they decay significantly in rainy and foggy weather [5], making LiDAR SLAM degrade in adverse weather conditions [4]. Noise commonly occurs in point cloud data due to hardware limitations, motion, lighting, dust, or reflective surface [6, 7, 8, 9, 10, 11]. Also, mechanical, solid-state, and phased-array LiDAR sensors are built differently and SLAM designed for one may not be interchangeable with others [4]. Instability causes localization to miscalculate a deviation from the agent’s current position. Depending on the scale of the deviation, it can cause automated vehicles to drive off-road or onto the wrong lane. In robotic systems, it would result in the wake-up or kidnapped robot problem, since the large gap between the two consecutive frames’ positions can cause the robot to be unable to recover, i.e., either it stops due to failed localization or reset and initiates a costly recovery and relocalize process, discarding the learned map/knowledge [12, 13, 14].
Such challenges to the LiDAR sensor raise concerns about the performance of SLAM in diverse environments [4]. To this end, there is a lack of comprehensive evaluation on the manifestation and impact of such issues on LiDAR SLAM. Moreover, increasing efforts invested in learning-based SLAM raises additional robustness concerns due to DNNs’ sensitivity to noises [15, 16, 17]. Yet there is a lack of comparative evaluation in that regard. Hence in this work, we evaluate the robustness of such applications against common LiDAR data corruption due to adverse weather, noisy environments, and losses. Once the effects are discovered, we want to explore whether current defense strategies such as re-training and denoising are sufficient to address these issues.
Our contributions can be summarized as follows:
  • We propose a novel framework to evaluate the robustness of the SOTA LiDAR-based localization against common LiDAR data corruptions. To the best of our knowledge, this is the first work to study this problem.
  • We study the effectiveness of common defense strategies, such as denoising and re-training, against common LiDAR data corruption.
  • We share our code and data publicly [^1] https://github.com/boyang9602/LiDARLocRobustness.
II Background and related works
LiDAR-based localization can either simultaneously build a map and localize the agent within it, i.e., SLAM [4], or localize the agent within a pre-built map, i.e., global localization [18].
II-A LiDAR-based localization and mapping
Figure 1: Different types of LiDAR SLAM Overview.
LiDAR-based localization is the most accurate, stable, resistant to different lighting conditions, and has no scale drift, thus suitable for large-scale applications [4]. It could be used from autonomous driving [19, 20], robot exploration [21] to deep space exploration [22]. LiDAR-based localization can be roughly categorized into LiDAR-only SLAM, MSF-based SLAM, and map-based global localization.
II-A1 LiDAR-only SLAM
LiDAR-only SLAM can be categorized into feature-based, direct, and projection-based methods as shown in Figure 1. Feature-based methods [23, 24] extract appearance, geometric, semantics, or shape-based features for scan matching. Alternatively, direct methods [25, 26] match the scan directly via handcrafted point cloud registration methods, neural networks, or a map. Finally, projection-based approaches [27, 28] convert the point cloud to a 2D representation and leverage existing visual SLAM methods. The effectiveness of these methods is highly dependent on the environments, equipment, and platforms. Traditional SLAM frameworks use hand-crafted features or models [29, 30, 31, 32, 33] to register scans. In contrast, learning-based SLAM [34, 35, 25], facilitated by the development of 3D computer vision, extracts features and register scans automatically. Due to the focus on computer vision, these methods only focus on feature extraction and scan registration rather than a complete solution. Moreover, their accuracy, stability, and performance leave much to be desired [4]. In our work, we look to evaluate the robustness of LiDAR-only SLAM.
II-A2 MSF-based SLAM
On the other hand, LiDAR can be incorporated into an MSF-based SLAM. MSF SLAM combines the output from the LiDAR with camera, IMU, and GNSS, to complement each other and improve localization performance [4, 36]. These methods build upon the LiDAR-only solutions. They are not included in this study.
II-A3 Map-based LiDAR Localization
Map-based localization registers an online LiDAR scan with a pre-built map and tries to locate it within the map [18, 37, 38, 39, 40, 41]. This offers precise and efficient localization, which is especially useful in safety-critical systems such as autonomous driving systems [42, 43]. However, these methods require a pre-built map, which is typically created using the SLAM algorithm offline and may require ground truth information to improve the map quality. Map-based localization shares many common technologies with LiDAR-only SLAM (e.g., LiDAR scan registration), so we will also attempt to evaluate the robustness of such systems.
II-B Robustness of LiDAR-based Systems
Previous works evaluated the impact of data corruption on LiDAR-based 3D obstacle detection. This includes benchmarking against perturbations [44, 45, 46, 47, 48]; testing and adversarial attacks via point-based [49, 50, 51, 52, 53, 54, 55, 56, 57] and object spoofing methods [58, 59, 60, 61, 62, 63, 64]. Finally, complex weather simulators [65, 66, 67, 68] model LiDAR’s interaction with dense droplets as well as how light bounces off objects under rain, snow, and fog. Many of these works are tailored to objects and thus inapplicable to localization. Hence, in this work, we evaluate LiDAR-based localization against a set of applicable perturbations.
III Methodology
In this section, we describe our framework for evaluating and improving the robustness of LiDAR-based localization.
Category Perturbation Type
Weather rain, snow, rain + wet ground (rain_wg),
snow + wet ground (snow_wg), fog
Noise background noise (bg_noise), upsample,
uniform noise in CCS (uni_noise),
gaussian noise in CCS (gau_noise),
impulse noise in CCS (imp_noise),
uniform noise in SCS (uni_noise_rad),
uniform noise in SCS (gau_noise_rad),
uniform noise in SCS (imp_noise_rad),
Density local density increase (local_inc),
local density decrease (local_dec),
beam deletion (beam_del),
layer deletion (layer_del), cutout
TABLE I: The 18 perturbations used by our work from three categories. CCS refers to Cartesian Coordinate System and SCS refers to Spherical Coordinate System. The details will be introduced in Section III.
III-A Evaluating the Robustness of LiDAR SLAM
LiDAR SLAMs are deployed in complex and diverse environments which may impact the sensitivity of LiDAR. It is important to evaluate the robustness and reliability of such systems in various environments. Our framework employs a total of 18 perturbations on cloud corruptions to emulate diverse environments such as adverse weather conditions and equipment malfunctions. We select a comprehensive set of corruption reflecting commonly occurring phenomena that impact the entire LiDAR scan instead of individual objects. Table I summarizes the list of perturbation types.
Weather Perturbations. Adverse weather such as rain and snow can negatively affect the quality of LiDAR scans. Randomly distributed droplets in the air from rain, snow, and fog can scatter light and disturb the measurement of LiDAR, thus injecting noises into the resulting PCD. Furthermore, wet ground always co-occurs with rain and snow due to the precipitation, which will reflect differently from dry ground.
Figure 2: An example of rain and wet ground corruption. The field view and reflection intensities are significantly decayed compared to the clean data.
Our framework simulates the following weather conditions considering both the direct impact of weather conditions on LiDAR (e.g., droplets causing scattering), as well as impacts of wet grounds on LiDAR (e.g., reflections).
  • Rain or Snow. We utilize an off-the-shelf weather simulator, i.e., LISA [65, 66] for simulating PCD under adverse weather such as rain and snow. LISA utilizes a physics-based model to simulate the decayed return of laser pulses in rain and snow weather, which is realistic as evaluated by previous work [47].
  • Fog. We utilize a fog Simulation on Real LiDAR Point Clouds (FS) [67] to simulate fog. While LISA already has a fog perturbation, it is more general, and works such as FS provide a better fog simulation model without needing unavailable data such as optical channel information. Hence, we use FS for fog simulation. FS simulates the effect of fog on real-world LiDAR point clouds that have been recorded in clear weather. FS models fog by modifying the impulse response changes to reflect the impact of different fog densities on the environment and the surrounding objects.
  • Rain + Wet Ground (rain_wg). We combine LISA and Wet Ground Model from [68] to produce a more realistic simulation than rain alone. The Wet Ground Model builds an optical model to simulate the reflection on the wet ground based on the thickness of the water. Figure 2 shows an example of this corruption.
  • Snow + Wet Ground (snow_wg). Similar to rain + wet ground, we produce more realistic snow weather by combining LISA’s snow simulation and the wet ground model on PCD.
Noise Perturbations. Noises happen during LiDAR measurement due to internal (e.g., equipment vibration and measurement errors) or external (e.g., dust and background light) factors [11, 69, 70]. We design perturbations on PCDs to simulate the errors attributed to such internal and external factors.
Noise Type Formula to Generate Noise
Gaussian (CCS) $p_{i}+\Delta p*SEV,\;\Delta p\sim\mathcal{N}(0,1)^{3},p_{i}\in P$
Uniform (CCS) $p_{i}+\Delta p*SEV,\;\Delta p\sim\mathcal{U}(-1,1)^{3},p_{i}\in P$
Impulse (CCS) $p_{i}+\Delta p*C,\;\Delta p\sim\mathcal{U}\{-1,1\}^{3},$
$p_{i}\in RS(P,SEV)$
Gaussian (SCS) $p_{i}+\Delta r*SEV,\;\Delta r\sim\mathcal{N}(0,1),p_{i}\in P$
Uniform (SCS) $p_{i}+\Delta r*SEV,\;\Delta r\sim\mathcal{U}(-1,1),p_{i}\in P$
Impulse (SCS) $p_{i}+\Delta r*C,\;\Delta r\sim\mathcal{U}\{-1,1\},$
$p_{i}\in RS(P,SEV)$
Background (CCS) $P_{new}=P\cup\{p_{i}^{\prime}|j=1,2,...,SEV\}$
$p_{i}^{\prime}=\{x_{i}^{\prime},y_{i}^{\prime},z_{i}^{\prime}\},\left\{\begin{array}[]{l}x_{i}^{\prime}\sim\mathcal{U}(x_{\min},x_{\max})\\ y_{i}^{\prime}\sim\mathcal{U}(y_{\min},y_{\max})\\ z_{i}^{\prime}\sim\mathcal{U}(z_{\min},z_{\max})\\ \end{array}\right.$
Upsample (CCS) $P_{new}=P\cup\{p_{i}+\Delta p|\Delta p\sim\mathcal{U}(-0.1,0.1),$
$p_{i}\in RS(P,SEV)\}$
TABLE II: Noise-based corruptions. SEV (severity) refers to the significance of corruption, whose definition varies based on the type of corruption. C refers to a constant value, whose concrete value varies in different places. RS refers to a random subset function, which samples a subset from the input based on the given $SEV$. $\Delta p$ refers to a point in CCS, so $+\Delta p$ means we change the three coordinates in CCS. $\Delta r$ refers to the range value, $+\Delta r$ means we only change the range value in SCS.
We show the details of noise perturbations in Table II. One frame of PCD is annotated as $P=\{p_{i}|i=1,2,...N\}$, where $N$ is the number of points in the point cloud. We apply three distributions of noises (Gaussian, Uniform, and Impulse) to two coordinate systems (Cartesian and Spherical, CCS and SCS for short) to emulate different noise sources. In CCS, one point is represented by $p=\{x,y,z\}$, i.e., Euclidean distance to each planes, while in SCS, one point is represented by $p=\{r,\phi,\theta\}$, where $r$ is the range, i.e., Euclidean distance from the origin to the point, $\phi$ is the azimuthal angle and $\theta$ is the polar angle. Noises in CCS represent inaccuracies caused by the vibration or rotating of the sensor [48] while SCS can better represent inaccurate measurement of time of flight (ToF) [47].
In addition to re-positioning existing PCD points, we synthesize new points to emulate environment noises, such as dust in the air, through background noise and upsample applied to CCS. Background noise adds points randomly in the entire space, while upsample adds points close to randomly selected points.
Density Perturbations.
Corruption Formula to Generate Noise
Local Inc $P_{new}=P\cup P_{add},$
$P_{add}=\bigcup_{i=0}^{SEV}IP(NN_{100}(p_{i}))$
Local Dec $P_{new}=P\setminus P_{rm},$
$P_{rm}=\bigcup_{i=0}^{SEV}RS(NN_{100}(p_{i}),C)$
Cutout $P_{new}=P\setminus\bigcup_{i=0}^{SEV}NN_{20}(p_{i})$
Beam Del $P_{new}=RS(P,SEV)$
Layer Del $P_{new}=P\setminus\bigcup_{i\in S}Layer_{i},$
$S=RS(\{1,2,...64\},SEV),P=\bigcup_{i=1}^{64}Layer_{i}$
TABLE III: Density-based corruptions. SEV (severity) refers to the significance of corruption, whose definition varies based on the type of corruption. C refers to a constant value, whose concrete value varies in different places. IP refers to the 3D interpolation function. $NN_{k}$ refers to the k-nearest neighbor function. RS refers to a random subset function, which samples a subset from the input based on the given number.
The density of points can be changed due to internal (e.g., malfunction of the equipment) or external (e.g., occlusion between objects and reflection on object surface) factors [71, 72]. We simulate five density-related corruptions, local density increase, local density decrease, cutout, beam deletion, and layer deletion. The definitions of these corruptions are listed in Table III.
III-B Improving the Robustness of LiDAR SLAM
Re-training. Learning-based systems are data-driven and could learn to fit themselves into a data distribution. We expect these systems to learn how to handle the corrupted data by themselves through re-training, which is a common enhancement and defensive strategy for deep learning [65, 67, 68, 73, 74].
Denoising. Handcraft algorithms are expected to be tuned and adapted by experts once inefficiencies are identified. Therefore such improvements are achieved case by case and lack of generality. Hence in this work, for non-learning-based SLAM, we focus on denoising the data, which is a more general way. Point cloud denoising removes unwanted noises from LiDAR scans to improve their quality. There are three categories of denoising, i.e., filter, optimization, and deep learning-based, each with its advantages and disadvantages [11]. In this work, we adopted a filter-based method, i.e., the bilateral filter by [75], which is reliable, and significantly faster than the alternatives while preserving the intensity with readily available implementation.
IV Experiment Setup
IV-A Subject LiDAR SLAM Systems
TABLE IV: Subject Systems. The errors are reported based on our replication.
Category Dataset Year Method Error
SLAM MULLS [76] 2021 Feature-based 0.56%
KISS-ICP [77] 2023 Direct 0.51%
DeLORA [78] 2021 Projection-based 7.78%
NeRF-LOAM [79] 2023 Distance Field 1.70%
Map-based
Localization LocNDF [37] 2023 Distance Field 0.059m
We experiment four SOTA LiDAR SLAM techniques and one map-based LiDAR localization, as shown in Table IV. We try to select one representative work from different types of mainstream LiDAR localization technologies, i.e. feature-based, direct, and projection-based methods [4]. We also include two neural distance field-based approaches, which have emerged and become popular since the success of the NeRF [80] in computer vision.
MULLS [76] is a feature-based SLAM that uses ground filtering and principal components analysis to extract features from the PCD. It builds local sub-maps using previous PCD and match the current frame with local sub-maps by the proposed multi-metric linear least square iterative closest point algorithm to compute the pose. KISS-ICP [77] is a direct SLAM that combines point-to-point ICP with adaptive thresholding, a robust kernel, universal motion compensation, and point cloud subsampling. DeLORA [78] is a projection-based, unsupervised SLAM. It projects the 3D point cloud to 2D image representations and employs ResNet to extract features from the 2D image. Then it predicts the translation and rotation by feeding the features into an MLP. NeRF-LOAM employs a neural network to predict the signed distance field (SDF). It first builds the previous frames into an octree. Then it samples intersected points between laser rays and voxels in the previously built octree. Given sampled points, the neural network predicts the SDF. Since sampled points are transformed to the same coordinate system as the previously built octree by the pose to be estimated, NeRF-LOAM optimizes the pose, neural network, and voxel embedding jointly by minimizing the loss between predicted SDF and real SDF. Different from NeRF-LOAM, LocNDF trains a neural network offline to predict the SDF. Given the online LiDAR scan, it tries to find a pose $P$ so that the neural network can produce minimal SDF when sampled points are transformed by $P$.
IV-B Datasets
We use KITTI Visual Odometry/SLAM Evaluation 2012 (KITTI) [19] to evaluate the four SLAM systems, DeLORA, MULLS, KISS-ICP, and NeRF-LOAM. KITTI 00 to 10 sequences are released with ground truths. We conducted our experiment on sequences 09 and 10 because (1) DeLORA was trained on sequences 00-08 hence only 09 and 10 remain. (2) sequences 09 and 10 are of moderate sizes which facilitate the large number of perturbations in a reasonable amount of time and computation resources.
Apollo South-Bay [81] contains completed ADS recordings covering various driving scenarios. In this work, we used a subset of ColumbiaPark-3, consisting of 700 testing frames, which is consistent with LocNDF’s pre-trained model [37].
IV-C Evaluation Metrics
For the evaluation of the precision of SLAMs, we adopted the relative pose error (RPE) [19]. RPE is widely used in previous work [4] to measure and compare the performance of SLAM. RPE is calculated as:
$RPE_{rot}(\mathscr{F})=\frac{1}{|\mathscr{F}|}\sum_{(i,j)\in\mathscr{F}}\angle[(\hat{p}_{j}\ominus\hat{p}_{i})\ominus(p_{j}\ominus p_{i})]$ (1)
$RPE_{trans}(\mathscr{F})=\frac{1}{|\mathscr{F}|}\sum_{(i,j)\in\mathscr{F}}||(\hat{p}_{j}\ominus\hat{p}_{i})\ominus(p_{j}\ominus p_{i})||_{2}$ (2)
where $RPE_{trans}$ measures the relative translation error and $RPE_{rot}$ measures the relative rotation error between the estimated pose and the ground truth, $\mathscr{F}$ is a set of frames $(i,j)$, $\hat{p}$ is the estimated pose and $p$ is the ground truth pose, $\ominus$ is the inverse compositional operator [82] and []delimited-[$\angle$[$\cdot$]∠ [ ⋅ ] is the rotation angle.
V Experiment Results
V-A Robustness Results of LiDAR SLAM
Method. We applied our robustness framework to evaluate the subject SLAM systems (Table IV). We first perturb the KITTI and Apollo-SouthBay datasets with the corruptions in Table I. Note that NeRF-LOAM is computationally expensive (i.e., requiring at least 24GB GPU memory and days to finish KITTI for one of the 18 perturbations), we took a statistically-significant sample of frames from KITTI for NeRF-LOAM as a proof of concept.
Figure 3: Performance of each system on corrupted KITTI LiDAR Odometry Sequence 09 and 10. Errors with significantly different magnitudes from the majority are aligned at their magnitude and the actual values are labeled in the figure. For example, an error of 12.79% is aligned to “$>$10.0”. If multiple errors overlap at the same magnitude, we label the two highest severity levels.
Results. We show the results of RPEtrans in Figures 3. Due to the space limitation, we do not present the RPErot in the paper as it is correlated with RPEtrans. We share the detailed results of RPErot in our public artifact.
For the non-learning SLAM systems, we find that they have degraded performance on almost all kinds of LiDAR perturbed data compared to the baseline (i.e., un-perturbed data). Furthermore, each SLAM system is vulnerable to a particular type of perturbation, where their performance drops significantly. The error of MULLS increases from below 0.5% to up to 0.8%. Even worse, the error of MULLS with “local density decrease” (severity 5) increased to more than 12% for KITTI 09 and 18% for KITTI 10 from 0.5%, which is more than 100 times worse. Similarly, kiss-icp is especially vulnerable to “background noise”. While the error only increased by less than 0.2% in most cases, the error of kiss-icp caused by background noise increases up to 80% (with severity 5 on KITTI 10) from less than 0.6% RPEtrans. Moreover, even the lightest background noise (severity 1) causes a significant error increase from less than 0.6% to more than 3% RPEtrans on KITTI 10.
We observe different results for the learning-based SLAM, Delora and NeRF-LOAM. Delora is vulnerable to most perturbations, especially those related to noise. The RPEtrans error increases from less than 10 percent to tens of percent (up to 50%). Lastly, NeRF-LOAM is quite different from others. We find the results differ very much between the two KITTI sequences, indicating that NeRF-LOAM’s robustness results highly depend on data. In KITTI 09, NeRF-LOAM is not sensitive to most perturbations except for one. While in KITTI 10, NeRF-LOAM is shown to be highly sensitive in most cases. Also, interestingly, in many cases, it shows better robustness with a higher severity than a lower severity. For example, for the cutout perturbation on KITTI 10, NeRF-LOAM performs poorly with severity 4, however, the performance is very well and much higher with severity 5 which is close to the baseline with the clean data. NeRF-LOAM is optimizing a mini-network online for tens of iterations. We hypothesize that the poor results are because NeRF-LOAM is far from being optimized in tens of iterations for these cases.
For the map-based localization system LocNDF, our evaluation shows that it is very robust against all types of corruptions. The map-based SLAM LocNDF is robust, at the cost of significant upfront investments (in terms of multiple scans of the environment to build the LiDAR map) and its inability to work in unexplored environments.
Our results show that SOTA Map-based localization, LocNDF, is robust against LiDAR data corruption under adverse weather and noises. The other SOTA SLAM systems including two learning-based SLAM are vulnerable to one or multiple types of LiDAR data corruption.
V-B Improving the Robustness of LiDAR SLAM
Method. We only concern MULLS, kiss-icp, and Delora in this denoising experiment, since LocNDF is robust against all kinds of corruption, while NeRF-LOAM could not produce consistent results between different trials on the same data. Firstly, we apply the Bilateral Filter to denoise the corrupted data and then re-run all the localization algorithms on the denoised data. Furthermore, since Delora uses a learning-based approach, we augment Delora’s training set (i.e., KITTI 00 to 08 sequences) by applying our proposed LiDAR corruptions and re-train Delora on the augmented data.
We focus on denoising the most representative corruption types that affect the algorithms most, i.e., local density decrease (MULLS and Delora), background noise (kiss-icp), uniform noise (Delora), Gaussian noise (Delora), and rain/snow (moderate effects on all systems). We denoise these types of corruption at severity 1, 3, and 5. We also denoise the clean data to see if the denoising has side effects on clean data.
Similarly, we fine-tune Delora on the corruptions that affect it most significantly, namely, rain, snow, fog, upsample, uniform noise, Gaussian noise, impulse noise, local density decrease, cutout, and layer deletion. We only augmented the training data with these corruptions at severity 5 and fine-tuned Delora for 6 epochs with the default training hyper-parameters.
Results.
Figure 4: (a) Performance on the denoised data. (b) Performance of the retrained model.
We present the performance of the three systems on the denoised data for KITTI 09 in Figure 4 (a). The results on KITTI 10 are available in our replication package due to the space limitation. For MULLS, local density decrease (local_dec) at severity 5 is the corruption that affects it most significantly. We find that the errors even increased by 2.43% after denoising. Besides, we find that denoising could reduce the error by 0.26% for Gaussian noise. The changes are insignificant for the remaining cases (less than 0.1%).
For kiss-icp, background noise (bg_noise) is the corruption that affects it most significantly. We find that the denoising can reduce the error by 0.2% (severity 1) to 3.22% (severity 5), that is, the negative effect of the corruption is fully eliminated (we observed the same effects on KITTI 10, where the errors at severity 5 are reduced by 80%). However, the denoising causes the errors on rain and snow to increase by 0.13% to 0.18%, and on local density decrease at severity 5 to increase by 0.28%. The changes are insignificant for the remaining cases.
For Delora, we find that denoising can reduce its errors in most cases including the baseline, by 0.44% to 16.08%. Similar to MULLS and kiss-icp, we find that the errors in local density decrease (local_dec) increased by 2.74%.
Bilateral Filter is beneficial to noise-based corruption, and can fully eliminate the negative effects of background noise. However, it does not help with density-based corruption (local density decrease).
We present the performance of the Delora after fine-turning for KITTI 09 in Figure 4 (b). We find that re-training can bring a much higher improvement than denoising: the errors are reduced by 2.73% to 35.79% for the baseline and all types of corruption. However, for KITTI 10, despite the overall effects being similar, we observed a slight increase in the errors for the baseline and a few corruptions. It is important to note that our fine-tuning was relatively naive: we only augmented KITTI 00 to 08 with the selected corruptions at severity 5 and used the original learning rate. We believe that further improvements could be achieved by a more refined fine-tuning process, such as augmenting the dataset with the corruption at different severities, and adjusting the learning rate and the ratio between clean and corrupted data.
Re-training with the corruption data is a more effective way to improve the robustness of the learning-based SLAM (Delora).
VI conclusion
In this work, we extensively evaluated the robustness of the LiDAR-based localization systems, including SLAM and map-based localization. Our results revealed that all kinds of SOTA LiDAR-based SLAM systems are vulnerable to specific types of LiDAR data corruption. We further showed that simple Bilateral Filter denoising can effectively eliminate noise-based data corruptions, however, the density-based corruptions remain unsolved. We also showed that the data augmentation with corrupted data can improve the robustness of the learning-based LiDAR SLAM significantly, and is even beneficial to its performance on clean data.
VII Acknowledgement
We would like to express our gratitude to our interns, Marko Griesser-Aleksic, Ali Ouabou El Ghoulb, and Hanchong Qiu, for their invaluable assistance and contribution to this project.

### v2

@@
-Evaluating and Improving the Robustness of LiDAR-based Localization and Mapping
+Robustness of LiDAR-Based Pose Estimation: Evaluating and Improving Odometry and Localization Under Common Point Cloud Corruptions
 Abstract
-LiDAR is one of the most commonly adopted sensors for simultaneous localization and mapping (SLAM) and map-based global localization.
-SLAM and map-based localization are crucial for the independent operation of autonomous systems, especially when external signals such as GNSS are unavailable or unreliable.
-While state-of-the-art (SOTA) LiDAR SLAM systems could achieve 0.5% (i.e., 0.5m per 100m) of errors and map-based localization could achieve centimeter-level global localization, it is still unclear how robust they are under various common LiDAR data corruptions.
-In this work, we extensively evaluated five SOTA LiDAR-based localization systems under 18 common scene-level LiDAR point cloud data (PCD) corruptions.
-We found that the robustness of LiDAR-based localization varies significantly depending on the category.
-For SLAM, hand-crafted methods are in general robust against most types of corruption, while being extremely vulnerable (up to +80% errors) to a specific corruption.
-Learning-based methods are vulnerable to most types of corruptions.
-For map-based global localization, we found that the SOTA is resistant to all applied corruptions.
-Finally, we found that simple Bilateral Filter denoising effectively eliminates noise-based corruption but is not helpful in density-based corruption.
-Re-training is more effective in defending learning-based SLAM against all types of corruption.
+Accurate and reliable pose estimation, i.e., determining the precise position and orientation of autonomous robots and vehicles, is critical for tasks like navigation and mapping.
+LiDAR is a widely used sensor for pose estimation, with odometry and localization being two primary tasks.
+LiDAR odometry estimates the relative motion between consecutive scans, while LiDAR localization aligns real-time scans with a pre-recorded map to obtain a global pose.
+Although they have different objectives and application scenarios, both rely on point cloud registration as the underlying technique and face shared challenges of data corruption caused by adverse conditions (e.g., rain).
+While state-of-the-art (SOTA) pose estimation systems achieved high accuracy on clean data, their robustness to corrupted data remains unclear.
+In this work, we propose a framework to systematically evaluate five SOTA LiDAR pose estimation systems across 18 synthetic real-world point cloud corruptions.
+Our experiments reveal that odometry systems degrade significantly under specific corruptions, with relative position errors increasing from 0.5% to more than 80%, while localization systems remain highly robust.
+We further demonstrate that denoising techniques can effectively mitigate the adverse effects of noise-induced corruptions, and re-training learning-based systems with corrupted data significantly enhances the robustness against various corruption types.
 I Introduction
-Robotics and automation systems have been gaining significant prevalence in this decade.
-Of those, robots, drones, and autonomous vehicles have been making the news [1] due to their widespread adoption and related safety concerns [2].
-Localization is fundamental to these systems since it keeps track of the position of a robotic system within an environment.
-Outdoors, and Global Navigation Satellite Systems (GNSS) are widely used for localization.
-In contrast, in urban, tunnel, and indoor settings where GNSS is hindered by weak satellite signals [3], simultaneous localization and mapping (SLAM) systems are favored since they operate independently of external signals and reliably navigate both indoors and outdoors.
-For these reasons, SLAM is widely used for locating and mapping-related tasks in robotics and autonomous driving.
-SLAM methods utilize either a camera, LiDAR, or a combination of both, with the optional integration of an Inertial Measurement Unit (IMU) [4].
-Compared to vision SLAM, LiDAR SLAM has higher mapping accuracy, and better stability, is less susceptible to lighting conditions, and does not observe scale drift.
-These advantages lead to its popularity in large-scale use cases and facilitate more research and application in this direction.
-Despite the recent successes of LiDAR SLAM, many challenges remain.
-Laser signals are robust under clear conditions, however, they decay significantly in rainy and foggy weather [5], making LiDAR SLAM degrade in adverse weather conditions [4].
-Noise commonly occurs in point cloud data due to hardware limitations, motion, lighting, dust, or reflective surface [6, 7, 8, 9, 10, 11].
-Also, mechanical, solid-state, and phased-array LiDAR sensors are built differently and SLAM designed for one may not be interchangeable with others [4].
-Instability causes localization to miscalculate a deviation from the agent’s current position.
-Depending on the scale of the deviation, it can cause automated vehicles to drive off-road or onto the wrong lane.
-In robotic systems, it would result in the wake-up or kidnapped robot problem, since the large gap between the two consecutive frames’ positions can cause the robot to be unable to recover, i.e., either it stops due to failed localization or reset and initiates a costly recovery and relocalize process, discarding the learned map/knowledge [12, 13, 14].
-Such challenges to the LiDAR sensor raise concerns about the performance of SLAM in diverse environments [4].
-To this end, there is a lack of comprehensive evaluation on the manifestation and impact of such issues on LiDAR SLAM.
-Moreover, increasing efforts invested in learning-based SLAM raises additional robustness concerns due to DNNs’ sensitivity to noises [15, 16, 17].
-Yet there is a lack of comparative evaluation in that regard.
-Hence in this work, we evaluate the robustness of such applications against common LiDAR data corruption due to adverse weather, noisy environments, and losses.
-Once the effects are discovered, we want to explore whether current defense strategies such as re-training and denoising are sufficient to address these issues.
-Our contributions can be summarized as follows:
-  • We propose a novel framework to evaluate the robustness of the SOTA LiDAR-based localization against common LiDAR data corruptions.
-To the best of our knowledge, this is the first work to study this problem.
-  • We study the effectiveness of common defense strategies, such as denoising and re-training, against common LiDAR data corruption.
-  • We share our code and data publicly [^1] https://github.com/boyang9602/LiDARLocRobustness.
+Autonomous robots and vehicles rely on real-time pose estimation to navigate, avoid collisions, and perform tasks such as mapping and path planning.
+LiDAR enables precise environmental perception by generating high-resolution 3D point clouds independent of external lighting conditions.
+As a result, it is widely used for pose estimation, particularly in scenarios where GNSS signals are unreliable (e.g., urban canyons) or unavailable (e.g., tunnels).
+Two primary methods are commonly employed for LiDAR-based pose estimation: odometry [1] and localization [2].
+LiDAR odometry estimates incremental motion by registering consecutive scans, making it essential for exploring unknown environments.
+However, odometry suffers from drift due to error accumulation over time.
+LiDAR localization, on the other hand, aligns online scans with pre-built maps to achieve drift-free global pose estimation.
+This is particularly crucial in safety-critical applications, such as fully autonomous vehicles that require global localization accuracy at the centimeter level [3].
+Despite the different objectives, both tasks share underlying point cloud registration techniques to obtain fine poses.
+Point cloud data corruptions can occur in various real-world scenarios, including adverse weather [4], noises from sensor [5] or dusts [6], reflective surfaces [7], and occlusions caused by dynamic environments [8, 9].
+Despite SOTA LiDAR-based odometry and localization methods achieving highly accurate results on clean LiDAR datasets, corrupted point cloud data will inevitably bring challenges to their performance [10].
+However, the extent to which their performance degrades under these corruptions remains unclear.
+Severe performance degradation could lead to critical failures, such as robots getting lost [11] and autonomous vehicles driving off-road [12].
+Though prior studies explored the robustness [9] and security [13] of LiDAR-based localization systems, there is still a lack of systematic evaluation of how point cloud corruptions impact LiDAR-based pose estimation systems.
+Our work bridges this gap with three key contributions:
+  • A framework to systematically evaluate the robustness of five SOTA LiDAR-based pose estimation systems under 18 common point cloud corruptions.
+  • An investigation of mitigation strategies, including denoising and re-training (for learning-based systems).
+  • Public release of code and data for reproducibility [^1] https://github.com/boyang9602/LiDARLocRobustness.
 II Background and related works
-LiDAR-based localization can either simultaneously build a map and localize the agent within it, i.e., SLAM [4], or localize the agent within a pre-built map, i.e., global localization [18].
-II-A LiDAR-based localization and mapping
-Figure 1: Different types of LiDAR SLAM Overview.
-LiDAR-based localization is the most accurate, stable, resistant to different lighting conditions, and has no scale drift, thus suitable for large-scale applications [4].
-It could be used from autonomous driving [19, 20], robot exploration [21] to deep space exploration [22].
-LiDAR-based localization can be roughly categorized into LiDAR-only SLAM, MSF-based SLAM, and map-based global localization.
-II-A1 LiDAR-only SLAM
-LiDAR-only SLAM can be categorized into feature-based, direct, and projection-based methods as shown in Figure 1.
-Feature-based methods [23, 24] extract appearance, geometric, semantics, or shape-based features for scan matching.
-Alternatively, direct methods [25, 26] match the scan directly via handcrafted point cloud registration methods, neural networks, or a map.
-Finally, projection-based approaches [27, 28] convert the point cloud to a 2D representation and leverage existing visual SLAM methods.
-The effectiveness of these methods is highly dependent on the environments, equipment, and platforms.
-Traditional SLAM frameworks use hand-crafted features or models [29, 30, 31, 32, 33] to register scans.
-In contrast, learning-based SLAM [34, 35, 25], facilitated by the development of 3D computer vision, extracts features and register scans automatically.
-Due to the focus on computer vision, these methods only focus on feature extraction and scan registration rather than a complete solution.
-Moreover, their accuracy, stability, and performance leave much to be desired [4].
-In our work, we look to evaluate the robustness of LiDAR-only SLAM.
-II-A2 MSF-based SLAM
-On the other hand, LiDAR can be incorporated into an MSF-based SLAM.
-MSF SLAM combines the output from the LiDAR with camera, IMU, and GNSS, to complement each other and improve localization performance [4, 36].
-These methods build upon the LiDAR-only solutions.
-They are not included in this study.
-II-A3 Map-based LiDAR Localization
-Map-based localization registers an online LiDAR scan with a pre-built map and tries to locate it within the map [18, 37, 38, 39, 40, 41].
-This offers precise and efficient localization, which is especially useful in safety-critical systems such as autonomous driving systems [42, 43].
-However, these methods require a pre-built map, which is typically created using the SLAM algorithm offline and may require ground truth information to improve the map quality.
-Map-based localization shares many common technologies with LiDAR-only SLAM (e.g., LiDAR scan registration), so we will also attempt to evaluate the robustness of such systems.
+II-A LiDAR-Based Pose Estimation
+LiDAR-based pose estimation allows autonomous agents to track their motions (odometry) and get their absolute position and orientation in a map (localization) in 3D environments.
+The core of LiDAR-based pose estimation is point cloud registration, i.e., aligning LiDAR scans to estimate their 6-degree-of-freedom (DoF) spatial transformations (translation and rotation).
+II-A1 Point Cloud Registration
+Given a source point cloud $\mathcal{P}\in\mathbb{R}^{3\times N}$ and a target point cloud $\mathcal{T}\in\mathbb{R}^{3\times M}$, point cloud registration solves:
+$\mathbf{R}^{*},\mathbf{t}^{*}=\underset{\mathbf{R}\in SO(3),\mathbf{t\in\mathbb{R}^{3}}}{\arg\min}\ \text{dist}\left(\mathbf{R}\mathcal{P}+\mathbf{t},\mathcal{T}\right)$ (1)
+Here, $\mathbf{R}$ is a 3-DoF rotation matrix, and $\mathbf{t}$ is a 3-DoF translation vector.
+The operation $\mathbf{R}\mathcal{P}+\mathbf{t}$ rotates all points using $\mathbf{R}$ and translates them by $\mathbf{t}$.
+The function $\text{dist}(\cdot,\cdot)$ quantifies the errors between the corresponding points in $\mathbf{R}\mathcal{P}+\mathbf{t}$ and $\mathcal{T}$.
+In other words, this equation seeks the optimal pose $\mathbf{R}^{*},\mathbf{t}^{*}$ that aligns the source point cloud $\mathcal{P}$ to the target point cloud $\mathcal{T}$.
+Various solutions have been proposed for point cloud registration, evolving from geometry-based methods to learning-based methods.
+They can be broadly categorized into direct matching, feature-based, and learning-based methods [1].
+Direct matching methods, such as Iterative Closest Point (ICP) [14], align raw points by iteratively minimizing the distance between nearest neighbor points.
+While vanilla ICP struggles with problems such as sensitivity to initialization and noise, modern variants such as KISS-ICP [15] address these limitations and remain competitive currently.
+Another direct method, Normal Distribution Transform (NDT) [16], partitions the target point cloud into a grid of cells, with each cell modeling the local distribution of points as a Gaussian.
+Then it finds the optimal pose that transforms the source point cloud to best fit the normal distribution in the target grid.
+Feature-based methods like LiDAR Odometry and Mapping (LOAM) [17] achieve fast and robust odometry by extracting and matching feature points such as sharp edges and planar surfaces.
+With the development of deep learning, researchers further improve feature extraction from hand-crafted features to learned features and predict the pose using neural networks [18].
+Instead of processing the 3D point cloud directly, [19] projects the points into 2D images and leverages the neural network to extract features from images and predict the pose.
+The success of Neural Radiance Fields (NeRF) [20] further inspires the recent exploration of implicit neural representations of scenes and advances the LiDAR-based pose estimations [21, 22].
+II-A2 LiDAR Odometry
+LiDAR odometry tracks the accurate motion of autonomous agents in real-time by aligning consecutive LiDAR scans.
+Despite the high accuracy of point cloud registration, the long-term trajectory drifts significantly due to accumulated errors.
+To mitigate the drift, additional algorithms such as loop closure [23] should be applied in post-processing.
+SOTA odometry systems cover all the approaches mentioned in the previous section [15, 17, 24, 19, 18, 22].
+II-A3 LiDAR Localization
+LiDAR localization determines the global 6-DoF pose by aligning an online scan with a pre-built map, which is critical for fully autonomous driving systems that require centimeter-level accuracy.
+While some localization systems estimate only 3-DoF position, we focus on 6-DoF pose estimation in this work.
+Compared to scan-to-scan registration in odometry, scan-to-map registration in localization is more challenging due to the vast search space.
+To address this problem, most methods first obtain an initial pose via external signals (e.g., GNSS) [3] or global position recognition [2], then perform the point cloud registration to estimate the fine pose.
+Although one-stage global pose estimation approaches exist, they generally underperform compared to those with initialization [2].
+Similar to odometry, SOTA localization approaches also cover the full spectrum of registration methods [3, 25, 26, 21].
+II-A4 Multi-Sensor Fusion Approaches
+Multi-sensor fusion (MSF) combines complementary sensors, such as LiDAR, cameras (vision odometry and localization), IMUs (high-frequency motion data), and GNSS (global localization), to enhance pose estimation [3].
+In this work, we focus on LiDAR-only pose estimation systems to study LiDAR’s standalone vulnerabilities in pose estimation.
 II-B Robustness of LiDAR-based Systems
-Previous works evaluated the impact of data corruption on LiDAR-based 3D obstacle detection.
-This includes benchmarking against perturbations [44, 45, 46, 47, 48]; testing and adversarial attacks via point-based [49, 50, 51, 52, 53, 54, 55, 56, 57] and object spoofing methods [58, 59, 60, 61, 62, 63, 64].
-Finally, complex weather simulators [65, 66, 67, 68] model LiDAR’s interaction with dense droplets as well as how light bounces off objects under rain, snow, and fog.
-Many of these works are tailored to objects and thus inapplicable to localization.
-Hence, in this work, we evaluate LiDAR-based localization against a set of applicable perturbations.
+Previous research has extensively explored the robustness and security of LiDAR-based 3D perception systems, particularly in obstacle detection.
+Several studies have proposed realistic adverse weather simulations for point cloud, including rain [27, 28], snow [29], and fog [30].
+Researchers also systematically benchmarked the LiDAR-based object detection on common point cloud data corruptions [31] and investigated its vulnerability to adversarial attacks [32].
+In the domain of LiDAR-based pose estimation, recent works have examined the security of localization and odometry against adversarial attacks [13, 33, 34].
+The most relevant study to ours is [9], which proposed a closed-form approximation of localization errors for the ICP algorithm.
+In contrast, our study systematically evaluates the robustness of multiple pose estimation systems, covering various underlying point cloud registration algorithms, against a comprehensive set of point cloud data corruptions.
 III Methodology
-In this section, we describe our framework for evaluating and improving the robustness of LiDAR-based localization.
+This section presents our framework for evaluating and improving the robustness of LiDAR-based pose estimation.
 Category Perturbation Type
@@
 The details will be introduced in Section III.
-III-A Evaluating the Robustness of LiDAR SLAM
-LiDAR SLAMs are deployed in complex and diverse environments which may impact the sensitivity of LiDAR.
-It is important to evaluate the robustness and reliability of such systems in various environments.
-Our framework employs a total of 18 perturbations on cloud corruptions to emulate diverse environments such as adverse weather conditions and equipment malfunctions.
-We select a comprehensive set of corruption reflecting commonly occurring phenomena that impact the entire LiDAR scan instead of individual objects.
-Table I summarizes the list of perturbation types.
+III-A Evaluating Robustness of LiDAR-Based Pose Estimation
+Autonomous robotics and vehicles are deployed in complex and diverse environments which affects the LiDAR performance and, consequently, the accuracy of pose estimation.
+To rigorously evaluate the robustness of pose estimation systems, our framework employs a broad spectrum of point cloud corruption (18 different types) that simulates adverse weather, internal and external noises, equipment malfunctions, and occlusions.
+Table I summarizes the list of corruption types.
 Weather Perturbations.
-Adverse weather such as rain and snow can negatively affect the quality of LiDAR scans.
-Randomly distributed droplets in the air from rain, snow, and fog can scatter light and disturb the measurement of LiDAR, thus injecting noises into the resulting PCD.
-Furthermore, wet ground always co-occurs with rain and snow due to the precipitation, which will reflect differently from dry ground.
-Figure 2: An example of rain and wet ground corruption. The field view and reflection intensities are significantly decayed compared to the clean data.
+Adverse weather can negatively affect the quality of LiDAR scans.
+Randomly distributed droplets in the air from rain, snow, and fog can scatter light and disturb the measurement of LiDAR, and cause corruptions in point cloud.
+Furthermore, wet ground always co-occurs with rain and snow due to the precipitation, which reflects differently from dry ground.
+Despite the availability of real-world datasets capturing various adverse weather conditions [35, 36], we opt for simulation methods on clean datasets for the following reasons:
+  • Realism. Existing simulation methods are physics model-based and highly resemble real-world phenomena [31, 27, 29, 30].
+  • Controllable.
+In real-world datasets, the severity of adverse weather is typically unknown and cannot be adjusted.
+Simulation, on the other hand, allows us to parameterize the intensity of these conditions.
+  • Sensor limitations. Real-world datasets are limited by their specific sensors, which may be incompatible with certain algorithms [10]. Simulation allows us to apply adverse weather on any dataset.
+  • Completeness. No dataset contains the full range of corruptions. Simulation allows us to apply all corruptions on a single dataset for a comprehensive and fair comparison.
+Figure 1: An example of rain and wet ground corruption.
+The field view and reflection intensities are significantly decayed compared to the clean data.
+The color bar indicates the intensities of the points.
 Our framework simulates the following weather conditions considering both the direct impact of weather conditions on LiDAR (e.g., droplets causing scattering), as well as impacts of wet grounds on LiDAR (e.g., reflections).
   • Rain or Snow.
-We utilize an off-the-shelf weather simulator, i.e., LISA [65, 66] for simulating PCD under adverse weather such as rain and snow.
-LISA utilizes a physics-based model to simulate the decayed return of laser pulses in rain and snow weather, which is realistic as evaluated by previous work [47].
+We utilize an off-the-shelf weather simulator, i.e., LISA [27, 28] to simulate rain and snow.
+LISA utilizes a physics model to simulate the decayed return of laser pulses in rain and snow weather.
   • Fog.
-We utilize a fog Simulation on Real LiDAR Point Clouds (FS) [67] to simulate fog.
-While LISA already has a fog perturbation, it is more general, and works such as FS provide a better fog simulation model without needing unavailable data such as optical channel information.
-Hence, we use FS for fog simulation.
-FS simulates the effect of fog on real-world LiDAR point clouds that have been recorded in clear weather.
+We utilize Fog Simulation (FS) [30] to simulate fog.
 FS models fog by modifying the impulse response changes to reflect the impact of different fog densities on the environment and the surrounding objects.
-  • Rain + Wet Ground (rain_wg).
-We combine LISA and Wet Ground Model from [68] to produce a more realistic simulation than rain alone.
+  • Rain or Snow + Wet Ground (rain/snow_wg).
+We combine LISA and Wet Ground Model from [29] to produce a more realistic simulation than rain/snow alone.
 The Wet Ground Model builds an optical model to simulate the reflection on the wet ground based on the thickness of the water.
 Figure 1 shows an example of this corruption.
-  • Snow + Wet Ground (snow_wg). Similar to rain + wet ground, we produce more realistic snow weather by combining LISA’s snow simulation and the wet ground model on PCD.
 Noise Perturbations.
-Noises happen during LiDAR measurement due to internal (e.g., equipment vibration and measurement errors) or external (e.g., dust and background light) factors [11, 69, 70].
-We design perturbations on PCDs to simulate the errors attributed to such internal and external factors.
+Noises occur in LiDAR measurement due to internal (e.g., equipment vibration) or external (e.g., dust) factors [37, 38].
+We design perturbations on point cloud to simulate the errors attributed to such factors.
 Noise Type Formula to Generate Noise
@@
 $p_{i}\in RS(P,SEV)\}$
-TABLE II: Noise-based corruptions.
+TABLE II: Noise-induced corruptions.
 SEV (severity) refers to the significance of corruption, whose definition varies based on the type of corruption.
@@
 We show the details of noise perturbations in Table II.
-One frame of PCD is annotated as $P=\{p_{i}|i=1,2,...N\}$, where $N$ is the number of points in the point cloud.
-We apply three distributions of noises (Gaussian, Uniform, and Impulse) to two coordinate systems (Cartesian and Spherical, CCS and SCS for short) to emulate different noise sources.
+One frame of point cloud is annotated as $\mathcal{P}=\{p_{i}|i=1,2,...N\}$, where $N$ is the number of points in the point cloud.
+We apply three distributions of noises (Gaussian, Uniform, and Impulse) to two coordinate systems (Cartesian and Spherical, CCS and SCS for short) to simulate different noise sources.
 In CCS, one point is represented by $p=\{x,y,z\}$, i.e., Euclidean distance to each planes, while in SCS, one point is represented by $p=\{r,\phi,\theta\}$, where $r$ is the range, i.e., Euclidean distance from the origin to the point, $\phi$ is the azimuthal angle and $\theta$ is the polar angle.
 Noises in CCS represent inaccuracies caused by the vibration or rotating of the sensor [39] while SCS can better represent inaccurate measurement of time of flight (ToF) [31].
-In addition to re-positioning existing PCD points, we synthesize new points to emulate environment noises, such as dust in the air, through background noise and upsample applied to CCS.
+In addition to re-positioning existing LiDAR points, we synthesize new points to simulate environment noises, such as dust in the air, through background noise and upsample applied to CCS.
 Background noise adds points randomly in the entire space, while upsample adds points close to randomly selected points.
@@
 $S=RS(\{1,2,...64\},SEV),P=\bigcup_{i=1}^{64}Layer_{i}$
-TABLE III: Density-based corruptions.
+TABLE III: Density-related corruptions.
 SEV (severity) refers to the significance of corruption, whose definition varies based on the type of corruption.
@@
 The definitions of these corruptions are listed in Table III.
-III-B Improving the Robustness of LiDAR SLAM
+III-B Improving Robustness of LiDAR-Based Pose Estimation
 Re-training.
 Learning-based systems are data-driven and could learn to fit themselves into a data distribution.
-We expect these systems to learn how to handle the corrupted data by themselves through re-training, which is a common enhancement and defensive strategy for deep learning [65, 67, 68, 73, 74].
+We expect these systems to learn how to handle the corrupted data by themselves through re-training, which is a common enhancement and defensive strategy for deep learning-based LiDAR-based algorithms [27, 29, 30].
 Denoising.
@@
 Therefore such improvements are achieved case by case and lack of generality.
-Hence in this work, for non-learning-based SLAM, we focus on denoising the data, which is a more general way.
+Hence in this work, for non-learning-based pose estimation systems, we focus on denoising the data, which is a more general way.
 Point cloud denoising removes unwanted noises from LiDAR scans to improve their quality.
@@
 IV Experiment Setup
-IV-A Subject LiDAR SLAM Systems
+IV-A Subject LiDAR-Based Pose Estimation Systems
 TABLE IV: Subject Systems. The errors are reported based on our replication.
 Category Dataset Year Method Error
-SLAM MULLS [76] 2021 Feature-based 0.56%
+Odometry MULLS [24] 2021 Feature-based 0.56%
 KISS-ICP [15] 2023 Direct 0.51%
-DeLORA [78] 2021 Projection-based 7.78%
-NeRF-LOAM [79] 2023 Distance Field 1.70%
-Map-based
-Localization LocNDF [37] 2023 Distance Field 0.059m
-We experiment four SOTA LiDAR SLAM techniques and one map-based LiDAR localization, as shown in Table IV.
-We try to select one representative work from different types of mainstream LiDAR localization technologies, i.e.
-feature-based, direct, and projection-based methods [4].
-We also include two neural distance field-based approaches, which have emerged and become popular since the success of the NeRF [80] in computer vision.
-MULLS [76] is a feature-based SLAM that uses ground filtering and principal components analysis to extract features from the PCD.
-It builds local sub-maps using previous PCD and match the current frame with local sub-maps by the proposed multi-metric linear least square iterative closest point algorithm to compute the pose.
-KISS-ICP [77] is a direct SLAM that combines point-to-point ICP with adaptive thresholding, a robust kernel, universal motion compensation, and point cloud subsampling.
-DeLORA [78] is a projection-based, unsupervised SLAM.
+Delora [19] 2021 Learning-based 7.78%
+NeRF-LOAM [22] 2023 NeRF-based 1.70%
+Localization LocNDF [21] 2023 NeRF-based 0.059m
+We experiment with four SOTA LiDAR odometry and one LiDAR localization systems, as shown in Table IV.
+We cover one representative work from different types of mainstream point cloud registration methods, i.e., feature-based, direct, and learning-based methods [10].
+We also include two neural distance field-based approaches, which have become popular since the success of the NeRF [20].
+MULLS [24] is a feature-based LiDAR odometry that uses ground filtering and principal components analysis to extract features from the point cloud.
+It builds local sub-maps using the previous point cloud and match the current frame with local sub-maps by the proposed multi-metric linear least square iterative closest point algorithm to compute the pose.
+KISS-ICP [15] is a direct matching LiDAR odometry that combines point-to-point ICP with adaptive thresholding, a robust kernel, universal motion compensation, and point cloud subsampling.
+Delora [19] is a learning-based, unsupervised LiDAR odometry.
 It projects the 3D point cloud to 2D image representations and employs ResNet to extract features from the 2D image.
-Then it predicts the translation and rotation by feeding the features into an MLP.
-NeRF-LOAM employs a neural network to predict the signed distance field (SDF).
+Then it predicts the translation and rotation by feeding the features into a neural network.
+NeRF-LOAM [22] employs a neural network to predict the signed distance field (SDF).
 It first builds the previous frames into an octree.
@@
 IV-B Datasets
-We use KITTI Visual Odometry/SLAM Evaluation 2012 (KITTI) [19] to evaluate the four SLAM systems, DeLORA, MULLS, KISS-ICP, and NeRF-LOAM.
+We use KITTI Visual Odometry/SLAM Evaluation 2012 (KITTI) [41] to evaluate the four odometry systems, Delora, MULLS, KISS-ICP, and NeRF-LOAM.
 KITTI 00 to 10 sequences are released with ground truths.
@@
 IV-C Evaluation Metrics
-For the evaluation of the precision of SLAMs, we adopted the relative pose error (RPE) [19].
-RPE is widely used in previous work [4] to measure and compare the performance of SLAM.
+For the evaluation of the precision of pose estimation, we adopted the relative pose error (RPE) [41].
+RPE is widely used in previous work [10] to measure and compare the performance of pose estimation.
 RPE is calculated as:
@@
 V Experiment Results
-V-A Robustness Results of LiDAR SLAM
+V-A Evaluating Robustness of LiDAR-Based Pose Estimation
 Method.
-We applied our robustness framework to evaluate the subject SLAM systems (Table IV).
+We applied our robustness framework to evaluate the subject odometry systems.
 We first perturb the KITTI and Apollo-SouthBay datasets with the corruptions in Table I.
-Note that NeRF-LOAM is computationally expensive (i.e., requiring at least 24GB GPU memory and days to finish KITTI for one of the 18 perturbations), we took a statistically-significant sample of frames from KITTI for NeRF-LOAM as a proof of concept.
+Note that NeRF-LOAM is computationally expensive (i.e., requiring at least 24GB GPU memory and days to finish one a single sequence), so we took a statistically-significant sample of frames from KITTI for NeRF-LOAM as a proof of concept.
 Figure 2: Performance of each system on corrupted KITTI LiDAR Odometry Sequence 09 and 10.
@@
 We share the detailed results of RPErot in our public artifact.
-For the non-learning SLAM systems, we find that they have degraded performance on almost all kinds of LiDAR perturbed data compared to the baseline (i.e., un-perturbed data).
-Furthermore, each SLAM system is vulnerable to a particular type of perturbation, where their performance drops significantly.
-The error of MULLS increases from below 0.5% to up to 0.8%.
-Even worse, the error of MULLS with “local density decrease” (severity 5) increased to more than 12% for KITTI 09 and 18% for KITTI 10 from 0.5%, which is more than 100 times worse.
+For the non-learning systems, we find that they have degraded performance on almost all kinds of LiDAR perturbed data compared to the baseline (i.e., un-perturbed data).
+Furthermore, each system is vulnerable to a particular type of perturbation, where their performance drops significantly.
+The error of MULLS increases from below 0.5% to up to 0.8% on most of the corruptions.
+However, the error of MULLS with “local density decrease” (severity 5) increased to more than 12% for KITTI 09 and 18% for KITTI 10.
 Similarly, kiss-icp is especially vulnerable to “background noise”.
@@
 Moreover, even the lightest background noise (severity 1) causes a significant error increase from less than 0.6% to more than 3% RPEtrans on KITTI 10.
-We observe different results for the learning-based SLAM, Delora and NeRF-LOAM.
+We observe different results for the learning-based systems, Delora and NeRF-LOAM.
 Delora is vulnerable to most perturbations, especially those related to noise.
@@
 Also, interestingly, in many cases, it shows better robustness with a higher severity than a lower severity.
-For example, for the cutout perturbation on KITTI 10, NeRF-LOAM performs poorly with severity 4, however, the performance is very well and much higher with severity 5 which is close to the baseline with the clean data.
-NeRF-LOAM is optimizing a mini-network online for tens of iterations.
-We hypothesize that the poor results are because NeRF-LOAM is far from being optimized in tens of iterations for these cases.
-For the map-based localization system LocNDF, our evaluation shows that it is very robust against all types of corruptions.
-The map-based SLAM LocNDF is robust, at the cost of significant upfront investments (in terms of multiple scans of the environment to build the LiDAR map) and its inability to work in unexplored environments.
-Our results show that SOTA Map-based localization, LocNDF, is robust against LiDAR data corruption under adverse weather and noises.
-The other SOTA SLAM systems including two learning-based SLAM are vulnerable to one or multiple types of LiDAR data corruption.
-V-B Improving the Robustness of LiDAR SLAM
+For example, for the cutout perturbation on KITTI 10, NeRF-LOAM performs poorly with severity 4, however, the performance is very well with severity 5 which is close to the baseline.
+NeRF-LOAM optimizes a mini-network online for tens of iterations.
+We hypothesize that the poor and unstable results are because NeRF-LOAM is far from being optimized in tens of iterations for these cases.
+For the LiDAR localization system LocNDF, our evaluation shows that it is very robust against all types of corruption.
+The robustness is at the cost of significant upfront investments (in terms of multiple scans of the environment to build the LiDAR map) and its inability to work in unexplored environments.
+Our results show that the SOTA LiDAR localization system, LocNDF, is robust against all point cloud corruptions.
+SOTA LiDAR odometry systems are vulnerable to one or multiple types of point cloud corruptions.
+V-B Improving Robustness of LiDAR-Based Pose Estimation
 Method.
 We only concern MULLS, kiss-icp, and Delora in this denoising experiment, since LocNDF is robust against all kinds of corruption, while NeRF-LOAM could not produce consistent results between different trials on the same data.
-Firstly, we apply the Bilateral Filter to denoise the corrupted data and then re-run all the localization algorithms on the denoised data.
-Furthermore, since Delora uses a learning-based approach, we augment Delora’s training set (i.e., KITTI 00 to 08 sequences) by applying our proposed LiDAR corruptions and re-train Delora on the augmented data.
+Firstly, we apply the bilateral filter to denoise the corrupted data and then re-run all the algorithms on the denoised data.
+Furthermore, since Delora uses a learning-based approach, we augment Delora’s training set (i.e., KITTI 00 to 08 sequences) by applying the proposed point cloud corruptions and re-train Delora on the augmented data.
 We focus on denoising the most representative corruption types that affect the algorithms most, i.e., local density decrease (MULLS and Delora), background noise (kiss-icp), uniform noise (Delora), Gaussian noise (Delora), and rain/snow (moderate effects on all systems).
@@
 We present the performance of the three systems on the denoised data for KITTI 09 in Figure 3 (a).
-The results on KITTI 10 are available in our replication package due to the space limitation.
+The results on KITTI 10 are similar and available in our replication package.
 For MULLS, local density decrease (local_dec) at severity 5 is the corruption that affects it most significantly.
@@
 Similar to MULLS and kiss-icp, we find that the errors in local density decrease (local_dec) increased by 2.74%.
-Bilateral Filter is beneficial to noise-based corruption, and can fully eliminate the negative effects of background noise.
-However, it does not help with density-based corruption (local density decrease).
+Bilateral filter can effectively mitigate noise-induced corruption in general, and fully eliminate the negative effects of background noise.
+However, it does not help with density-related corruption (local density decrease).
 We present the performance of the Delora after fine-turning for KITTI 09 in Figure 3 (b).
@@
 We believe that further improvements could be achieved by a more refined fine-tuning process, such as augmenting the dataset with the corruption at different severities, and adjusting the learning rate and the ratio between clean and corrupted data.
-Re-training with the corruption data is a more effective way to improve the robustness of the learning-based SLAM (Delora).
+Re-training with the corruption data is a more effective way to improve the robustness of the learning-based LiDAR pose estimation system (Delora).
 VI conclusion
-In this work, we extensively evaluated the robustness of the LiDAR-based localization systems, including SLAM and map-based localization.
-Our results revealed that all kinds of SOTA LiDAR-based SLAM systems are vulnerable to specific types of LiDAR data corruption.
-We further showed that simple Bilateral Filter denoising can effectively eliminate noise-based data corruptions, however, the density-based corruptions remain unsolved.
-We also showed that the data augmentation with corrupted data can improve the robustness of the learning-based LiDAR SLAM significantly, and is even beneficial to its performance on clean data.
-VII Acknowledgement
-We would like to express our gratitude to our interns, Marko Griesser-Aleksic, Ali Ouabou El Ghoulb, and Hanchong Qiu, for their invaluable assistance and contribution to this project.
+In this work, we extensively evaluated the robustness of LiDAR-based pose estimation systems, including odometry and localization, against 18 synthetic real-world point cloud corruptions.
+Our findings indicate that all SOTA LiDAR-based odometry systems are particularly vulnerable to certain corruption types.
+We further show that simple bilateral filtering effectively mitigates noise-induced data corruptions, however, density-related corruptions remain challenging.
+Finally, we show that corruption-based data augmentation not only significantly enhances the robustness of learning-based LiDAR odometry systems but also improves their performance on clean data.

### v3

@@
-Robustness of LiDAR-Based Pose Estimation: Evaluating and Improving Odometry and Localization Under Common Point Cloud Corruptions
+Evaluating and Improving the Robustness of LiDAR Odometry and Localization Under Real-World Corruptions
 Abstract
-Accurate and reliable pose estimation, i.e., determining the precise position and orientation of autonomous robots and vehicles, is critical for tasks like navigation and mapping.
-LiDAR is a widely used sensor for pose estimation, with odometry and localization being two primary tasks.
-LiDAR odometry estimates the relative motion between consecutive scans, while LiDAR localization aligns real-time scans with a pre-recorded map to obtain a global pose.
-Although they have different objectives and application scenarios, both rely on point cloud registration as the underlying technique and face shared challenges of data corruption caused by adverse conditions (e.g., rain).
-While state-of-the-art (SOTA) pose estimation systems achieved high accuracy on clean data, their robustness to corrupted data remains unclear.
-In this work, we propose a framework to systematically evaluate five SOTA LiDAR pose estimation systems across 18 synthetic real-world point cloud corruptions.
-Our experiments reveal that odometry systems degrade significantly under specific corruptions, with relative position errors increasing from 0.5% to more than 80%, while localization systems remain highly robust.
-We further demonstrate that denoising techniques can effectively mitigate the adverse effects of noise-induced corruptions, and re-training learning-based systems with corrupted data significantly enhances the robustness against various corruption types.
+LiDAR odometry and localization are two widely used and fundamental applications in robotic and autonomous driving systems.
+Although state‑of‑the‑art (SOTA) systems achieve high accuracy on clean point clouds, their robustness to corrupted data remains largely unexplored.
+We present the first comprehensive benchmark to evaluate the robustness of LiDAR pose‑estimation techniques under 18 realistic synthetic corruptions.
+Our results show that, under these corruptions, odometry position errors escalate from 0.5% to more than 80%, while localization performance stays consistently high.
+To address this sensitivity, we propose two complementary strategies.
+First, we design a lightweight detection-and-filter pipeline that classifies the point cloud corruption and applies a corresponding filter (e.g., bilateral filter for noise) to restore the point cloud quality.
+Our classifier accurately identifies each corruption type, and the filter effectively restores odometry accuracy to near-clean data levels.
+Second, for learning‑based systems, we show that fine-tuning using the corrupted data substantially improves robustness across all tested corruptions and even boosts performance on clean point clouds on one data sequence.
 I Introduction
-Autonomous robots and vehicles rely on real-time pose estimation to navigate, avoid collisions, and perform tasks such as mapping and path planning.
-LiDAR enables precise environmental perception by generating high-resolution 3D point clouds independent of external lighting conditions.
-As a result, it is widely used for pose estimation, particularly in scenarios where GNSS signals are unreliable (e.g., urban canyons) or unavailable (e.g., tunnels).
-Two primary methods are commonly employed for LiDAR-based pose estimation: odometry [1] and localization [2].
-LiDAR odometry estimates incremental motion by registering consecutive scans, making it essential for exploring unknown environments.
-However, odometry suffers from drift due to error accumulation over time.
-LiDAR localization, on the other hand, aligns online scans with pre-built maps to achieve drift-free global pose estimation.
-This is particularly crucial in safety-critical applications, such as fully autonomous vehicles that require global localization accuracy at the centimeter level [3].
-Despite the different objectives, both tasks share underlying point cloud registration techniques to obtain fine poses.
-Point cloud data corruptions can occur in various real-world scenarios, including adverse weather [4], noises from sensor [5] or dusts [6], reflective surfaces [7], and occlusions caused by dynamic environments [8, 9].
-Despite SOTA LiDAR-based odometry and localization methods achieving highly accurate results on clean LiDAR datasets, corrupted point cloud data will inevitably bring challenges to their performance [10].
-However, the extent to which their performance degrades under these corruptions remains unclear.
-Severe performance degradation could lead to critical failures, such as robots getting lost [11] and autonomous vehicles driving off-road [12].
-Though prior studies explored the robustness [9] and security [13] of LiDAR-based localization systems, there is still a lack of systematic evaluation of how point cloud corruptions impact LiDAR-based pose estimation systems.
-Our work bridges this gap with three key contributions:
-  • A framework to systematically evaluate the robustness of five SOTA LiDAR-based pose estimation systems under 18 common point cloud corruptions.
-  • An investigation of mitigation strategies, including denoising and re-training (for learning-based systems).
-  • Public release of code and data for reproducibility [^1] https://github.com/boyang9602/LiDARLocRobustness.
+Reliable pose estimation is critical in robotics and autonomous driving systems (ADS), where even minor errors can lead to catastrophic failures in navigation, planning, and control [1, 2].
+LiDAR-based pose estimation is widely adopted for its superior reliability and performance, particularly when Global Navigation Satellite System (GNSS) signals are unreliable (e.g., urban canyons) or unavailable (e.g., indoors).
+To this end, extensive research has focused on improving LiDAR pose estimation, i.e., odometry and localization.
+Particularly, LiDAR odometry estimates an agent’s incremental motion between consecutive LiDAR scans, enabling reliable self-tracking and map building in unknown environments.
+In contrast, LiDAR localization determines an agent’s absolute pose within a pre-built map, supporting accurate navigation in structured settings such as urban road networks.
+These methods vary in their registration strategies, including direct scan alignment [3, 4], feature-based matching [5], and, more recently, learning-based approaches [6, 7] to improve generalization across diverse environments.
+Despite the success, LiDAR odometry and localization face inherent challenges on robustness caused by non-clean LiDAR data due to real-world corruptions.
+Prior works reveal that point cloud data (PCD) may be corrupted or noisy, caused by real-world environment challenges, such as adverse weather [8], sensor noises [9] or dusts [10], reflective surfaces [11], and occlusions caused by dynamic environments [12].
+This issue is particularly pronounced in learning-based approaches, which often lack transparency and interpretability, making it difficult to understand their decision-making processes and limiting their ability to generalize to scenarios that diverge from the training data [13].
+While prior studies have examined specific aspects of robustness, e.g., occlusion in  [12] and security threats via adversarial attacks  [14], these works are limited in scope and do not offer a comprehensive robustness benchmark for LiDAR pose estimation systems.
+Notably, there is no systematic evaluation of how a wide range of real-world point cloud corruptions affect LiDAR odometry and localization, particularly in learning-based approaches, which are known to be more sensitive to data quality.
+Furthermore, existing efforts fall short in proposing methods to improve the robustness of LiDAR pose estimation against such corruptions.
+In this work, we propose a systematic evaluation framework, RobustLOL, to address the research gap regarding the impact of PCD corruptions on LiDAR odometry and localization, especially learning-based methods which are known to be sensitive to such corruptions.
+Furthermore, to mitigate these adverse effects, we further introduce and experiment with two complementary strategies to enhance robustness.
+First, we design a novel and lightweight detection‑and‑filter pipeline and validate its effectiveness on noise‑induced corruptions.
+Second, for learning‑based methods, we show that fine-tuning on corrupted data significantly enhances robustness.
+Our key contributions are as follows:
+  • We propose RobustLOL, a novel comprehensive robustness benchmark for LiDAR odometry/localization systems.
+  • We propose a novel lightweight detection-and-filter method which enables accurate, efficient, generalizable classification and restoration of corrupted PCD.
+  • We investigate various mitigation strategies, including detection-and-filter and fine-tuning to defend the LiDAR odometry/localization systems against PCD corruptions.
+  • Public release of code and data for reproducibility [^1] https://doi.org/10.5281/zenodo.17128983.
+Our framework reveals that SOTA LiDAR odometry is sensitive to at least one type of PCD corruptions.
+Our detection-and-filter can effectively defend the noise-induced corruptions for non-learning systems and corruption augmented fine-tuning can bring learning-based systems to near-clean level.
 II Background and Related Works
-II-A LiDAR-Based Pose Estimation
-LiDAR-based pose estimation allows autonomous agents to track their motions (odometry) and get their absolute position and orientation in a map (localization) in 3D environments.
-The core of LiDAR-based pose estimation is point cloud registration, i.e., aligning LiDAR scans to estimate their 6-degree-of-freedom (DoF) spatial transformations (translation and rotation).
+II-A LiDAR-Based Odometry and Localization
+LiDAR-based odometry and localization allow autonomous agents to track their motions (odometry) and get their absolute position and orientation in a map (localization) in 3D environments.
+The core of LiDAR-based odometry and localization is point cloud registration, the process of aligning two LiDAR scans to determine their relative poses.
 II-A1 Point Cloud Registration
@@
 $\mathbf{R}^{*},\mathbf{t}^{*}=\underset{\mathbf{R}\in SO(3),\mathbf{t\in\mathbb{R}^{3}}}{\arg\min}\ \text{dist}\left(\mathbf{R}\mathcal{P}+\mathbf{t},\mathcal{T}\right)$ (1)
-Here, $\mathbf{R}$ is a 3-DoF rotation matrix, and $\mathbf{t}$ is a 3-DoF translation vector.
-The operation $\mathbf{R}\mathcal{P}+\mathbf{t}$ rotates all points using $\mathbf{R}$ and translates them by $\mathbf{t}$.
+where $\mathbf{R}$ is a 3-DoF rotation matrix and $\mathbf{t}$ is a 3-DoF translation vector.
+The source point cloud is rotated by $\mathbf{R}$ and translated by $\mathbf{t}$.
 The function $\text{dist}(\cdot,\cdot)$ quantifies the errors between the corresponding points in $\mathbf{R}\mathcal{P}+\mathbf{t}$ and $\mathcal{T}$.
 In other words, this equation seeks the optimal pose $\mathbf{R}^{*},\mathbf{t}^{*}$ that aligns the source point cloud $\mathcal{P}$ to the target point cloud $\mathcal{T}$.
-Various solutions have been proposed for point cloud registration, evolving from geometry-based methods to learning-based methods.
-They can be broadly categorized into direct matching, feature-based, and learning-based methods [1].
-Direct matching methods, such as Iterative Closest Point (ICP) [14], align raw points by iteratively minimizing the distance between nearest neighbor points.
-While vanilla ICP struggles with problems such as sensitivity to initialization and noise, modern variants such as KISS-ICP [15] address these limitations and remain competitive currently.
-Another direct method, Normal Distribution Transform (NDT) [16], partitions the target point cloud into a grid of cells, with each cell modeling the local distribution of points as a Gaussian.
-Then it finds the optimal pose that transforms the source point cloud to best fit the normal distribution in the target grid.
-Feature-based methods like LiDAR Odometry and Mapping (LOAM) [17] achieve fast and robust odometry by extracting and matching feature points such as sharp edges and planar surfaces.
-With the development of deep learning, researchers further improve feature extraction from hand-crafted features to learned features and predict the pose using neural networks [18].
-Instead of processing the 3D point cloud directly, [19] projects the points into 2D images and leverages the neural network to extract features from images and predict the pose.
-The success of Neural Radiance Fields (NeRF) [20] further inspires the recent exploration of implicit neural representations of scenes and advances the LiDAR-based pose estimations [21, 22].
+Various approaches are proposed to solve this problem, evolving from geometry-based matching methods to modern learning-based methods.
+They can be broadly categorized into direct matching, feature-based, projection-based, and neural distance field-based methods.
+Direct methods like iterative closest point (ICP) [15, 4] align raw points by iteratively minimizing the distance between nearest neighbor points.
+Normal distribution transform (NDT) is another classic direct method [3], which partitions the target point cloud into a grid of cells, with each cell modeling the local distribution of points as a Gaussian.
+Instead of minimizing the distances between points, it optimizes the likelihood measure that evaluates how well the source point cloud fits the normal distribution in the target grid.
+A recent breakthrough is LiDAR Odometry and Mapping (LOAM) [5] which advances the point cloud registration by extracting robust features (e.g., edges and planar regions) from point clouds and matching the features.
+With the development of deep learning, researchers further improved feature extraction from hand-crafted features to learned features [6, 7].
+The success of Neural Radiance Fields (NeRF) [16] in modeling implicit 3D scenes has inspired the efforts to explore of the application in the LiDAR-based pose estimation [17, 18].
 II-A2 LiDAR Odometry
-LiDAR odometry tracks the accurate motion of autonomous agents in real-time by aligning consecutive LiDAR scans.
-Despite the high accuracy of point cloud registration, the long-term trajectory drifts significantly due to accumulated errors.
-To mitigate the drift, additional algorithms such as loop closure [23] should be applied in post-processing.
-SOTA odometry systems cover all the approaches mentioned in the previous section [15, 17, 24, 19, 18, 22].
+LiDAR odometry tracks real-time motion by aligning consecutive scans.
+While point cloud registration offers high accuracy, accumulated errors lead to significant long-term drift.
+To correct this, post-processing techniques like loop closure [19] are often used.
+SOTA odometry systems cover the various approaches as discussed above [4, 5, 20, 7, 6, 18].
 II-A3 LiDAR Localization
-LiDAR localization determines the global 6-DoF pose by aligning an online scan with a pre-built map, which is critical for fully autonomous driving systems that require centimeter-level accuracy.
-While some localization systems estimate only 3-DoF position, we focus on 6-DoF pose estimation in this work.
+LiDAR localization determines the global 6-DoF pose by aligning an online scan with a pre-built map, enabling centimeter-level accuracy critical for fully autonomous driving systems.
+While some localization systems estimate only 3-DoF position, our work focus on 6-DoF pose estimation.
 Compared to scan-to-scan registration in odometry, scan-to-map registration in localization is more challenging due to the vast search space.
@@
 Multi-sensor fusion (MSF) combines complementary sensors, such as LiDAR, cameras (vision odometry and localization), IMUs (high-frequency motion data), and GNSS (global localization), to enhance pose estimation [21].
-In this work, we focus on LiDAR-only pose estimation systems to study LiDAR’s standalone vulnerabilities in pose estimation.
+In this work, we focus on LiDAR-only pose estimation systems to study LiDAR’s standalone sensitivity to noises in pose estimation.
 II-B Robustness of LiDAR-based Systems
-Previous research has extensively explored the robustness and security of LiDAR-based 3D perception systems, particularly in obstacle detection.
-Several studies have proposed realistic adverse weather simulations for point cloud, including rain [27, 28], snow [29], and fog [30].
-Researchers also systematically benchmarked the LiDAR-based object detection on common point cloud data corruptions [31] and investigated its vulnerability to adversarial attacks [32].
-In the domain of LiDAR-based pose estimation, recent works have examined the security of localization and odometry against adversarial attacks [13, 33, 34].
-The most relevant study to ours is [9], which proposed a closed-form approximation of localization errors for the ICP algorithm.
-In contrast, our study systematically evaluates the robustness of multiple pose estimation systems, covering various underlying point cloud registration algorithms, against a comprehensive set of point cloud data corruptions.
+Previous research has extensively explored the robustness and security of LiDAR-based 3D obstacle detection.
+Several studies have also introduced realistic adverse weather simulations to evaluate the impact of environmental conditions on PCD, including rain [25, 26], snow [27], and fog [28].
+Researchers have also systematically benchmarked 3D obstacle detection under common point cloud corruptions [29] and investigated its vulnerability to adversarial attacks [30, 31].
+For LiDAR pose estimation, recent work has investigated the security of localization and odometry against adversarial attacks [14, 32, 33].
+The study most closely related to ours is Laconte et al.
+[12], which derives a closed-form approximation of the worst-case localization error for the ICP algorithm [15] as a function of the number of corrupted points.
+Their theoretical bound offers useful insight into how adversarial outliers can degrade a linearized (vanilla) ICP.
+However, SOTA odometry and localization systems typically employ more advanced ICP variants or alternative registration approaches, and the idealized bounds in [12] do not fully reflect the degradation under realistic corruptions.
+Our work complements this research gap by performing a systematic empirical evaluation of a broad set of odometry and localization pipelines, including both learning-based and classical registration methods, under a comprehensive suite of realistic corruption scenarios.
 III Methodology
-This section presents our framework for evaluating and improving the robustness of LiDAR-based pose estimation.
+In this section, we introduce a unified framework to evaluate LiDAR-based odometry and localization systems under common data corruptions, and we describe two defense strategies: a detection-and-filter pipeline and model fine-tuning on augmented data.
 Category Perturbation Type
-Weather rain, snow, rain + wet ground (rain_wg),
+Weather (5) rain, snow, rain + wet ground (rain_wg),
 snow + wet ground (snow_wg), fog
-Noise background noise (bg_noise), upsample,
+Noise (8) background noise (bg_noise), upsample,
 uniform noise in CCS (uni_noise),
@@
 uniform noise in SCS (uni_noise_rad),
-uniform noise in SCS (gau_noise_rad),
-uniform noise in SCS (imp_noise_rad),
-Density local density increase (local_inc),
+gaussian noise in SCS (gau_noise_rad),
+impulse noise in SCS (imp_noise_rad),
+Density (5) local density increase (local_inc),
 local density decrease (local_dec),
@@
 layer deletion (layer_del), cutout
-TABLE I: The 18 perturbations used by our work from three categories.
+TABLE I: Our robustness framework consists of 18 real-world LiDAR corruptions.
 CCS refers to Cartesian Coordinate System and SCS refers to Spherical Coordinate System.
-The details will be introduced in Section III.
-III-A Evaluating Robustness of LiDAR-Based Pose Estimation
-Autonomous robotics and vehicles are deployed in complex and diverse environments which affects the LiDAR performance and, consequently, the accuracy of pose estimation.
-To rigorously evaluate the robustness of pose estimation systems, our framework employs a broad spectrum of point cloud corruption (18 different types) that simulates adverse weather, internal and external noises, equipment malfunctions, and occlusions.
-Table I summarizes the list of corruption types.
+Their calculation formulas are listed in Table II and III.
+III-A RobustLOL: A Robustness Framework for LiDAR Odometry and Localization
+Autonomous robots and vehicles are deployed in complex and diverse environments.
+The performance of LiDAR may degrade under adverse conditions.
+To rigorously evaluate the robustness of odometry and localization systems, our framework employs 18 different types of PCD corruption (Table I) that simulate adverse weather, internal and external noises, equipment malfunctions, and occlusions.
 Weather Perturbations.
 Adverse weather can negatively affect the quality of LiDAR scans.
-Randomly distributed droplets in the air from rain, snow, and fog can scatter light and disturb the measurement of LiDAR, and cause corruptions in point cloud.
-Furthermore, wet ground always co-occurs with rain and snow due to the precipitation, which reflects differently from dry ground.
-Despite the availability of real-world datasets capturing various adverse weather conditions [35, 36], we opt for simulation methods on clean datasets for the following reasons:
-  • Realism. Existing simulation methods are physics model-based and highly resemble real-world phenomena [31, 27, 29, 30].
-  • Controllable.
-In real-world datasets, the severity of adverse weather is typically unknown and cannot be adjusted.
-Simulation, on the other hand, allows us to parameterize the intensity of these conditions.
-  • Sensor limitations. Real-world datasets are limited by their specific sensors, which may be incompatible with certain algorithms [10]. Simulation allows us to apply adverse weather on any dataset.
-  • Completeness. No dataset contains the full range of corruptions. Simulation allows us to apply all corruptions on a single dataset for a comprehensive and fair comparison.
+Randomly distributed airborne droplets from rain, snow, and fog scatter light, corrupting the PCD captured by LiDAR.
+Furthermore, precipitation results in wet ground that always co-occur with rain or snow, leading to surface reflections that differ significantly from dry ground.
+Despite the availability of real-world datasets capturing various adverse weather conditions [34, 35], we opt for simulation using clean datasets for three reasons.
+First, real datasets does not encompass the full spectrum of point-cloud corruptions, whereas simulation enables us to apply every corruptions on the same dataset, ensuring a comprehensive and fair comparison.
+Second, real recordings are sensor-specific, which can restrict compatibility with certain odometry or localization algorithms [13]; simulation, by contrast, is sensor-agnostic and can be applied to any dataset.
+Finally, the severity of weather effects in real recordings is fixed and often undocumented, but simulation lets us precisely parameterize corruption intensity and systematically study its impact, leveraging realistic models of fog, rain, and snow [29, 25, 27, 28].
 Figure 1: An example of rain and wet ground corruption.
 The field view and reflection intensities are significantly decayed compared to the clean data.
-The color bar indicates the intensities of the points.
-Our framework simulates the following weather conditions considering both the direct impact of weather conditions on LiDAR (e.g., droplets causing scattering), as well as impacts of wet grounds on LiDAR (e.g., reflections).
-  • Rain or Snow.
-We utilize an off-the-shelf weather simulator, i.e., LISA [27, 28] to simulate rain and snow.
-LISA utilizes a physics model to simulate the decayed return of laser pulses in rain and snow weather.
-  • Fog.
-We utilize Fog Simulation (FS) [30] to simulate fog.
-FS models fog by modifying the impulse response changes to reflect the impact of different fog densities on the environment and the surrounding objects.
+The numbers above the color bar show the intensities of the points with corresponding colors.
+Our framework simulates the following weather conditions considering both the direct impact of weather on LiDAR (e.g., droplets causing scattering), as well as impacts of wet grounds on LiDAR (e.g., reflections).
+  • Rain or Snow. We use LISA [25, 26], an off-the-shelf rain and snow simulator which use a physics model to simulate the decayed return of laser pulses during rain and snow.
+  • Fog. Following previous work [28], we model fog by altering the impulse response, capturing the effect of varying fog densities on the surrounding objects and environment.
   • Rain or Snow + Wet Ground (rain/snow_wg).
-We combine LISA and Wet Ground Model from [29] to produce a more realistic simulation than rain/snow alone.
-The Wet Ground Model builds an optical model to simulate the reflection on the wet ground based on the thickness of the water.
-Figure 1 shows an example of this corruption.
+We combine LISA and Wet Ground Model (WGM) [27] to produce a more realistic simulation than rain/snow alone.
+WGM builds an optical model to simulate wet ground reflection based on water thickness.
+We show an example of this corruption in Figure 1.
 Noise Perturbations.
-Noises occur in LiDAR measurement due to internal (e.g., equipment vibration) or external (e.g., dust) factors [37, 38].
+Noises occur in LiDAR measurement due to internal (e.g., equipment vibration) or external (e.g., dust or strong light) factors [36, 37].
 We design perturbations on point cloud to simulate the errors attributed to such factors.
@@
 $p_{i}\in RS(P,SEV)$
-Background (CCS) $P_{new}=P\cup\{p_{i}^{\prime}|j=1,2,...,SEV\}$
-$p_{i}^{\prime}=\{x_{i}^{\prime},y_{i}^{\prime},z_{i}^{\prime}\},\left\{\begin{array}[]{l}x_{i}^{\prime}\sim\mathcal{U}(x_{\min},x_{\max})\\ y_{i}^{\prime}\sim\mathcal{U}(y_{\min},y_{\max})\\ z_{i}^{\prime}\sim\mathcal{U}(z_{\min},z_{\max})\\ \end{array}\right.$
-Upsample (CCS) $P_{new}=P\cup\{p_{i}+\Delta p|\Delta p\sim\mathcal{U}(-0.1,0.1),$
-$p_{i}\in RS(P,SEV)\}$
+Background $P_{new}=P\cup\{p_{i}^{\prime}|j=1,2,...,SEV\}$
+[…]
+Upsample $P_{new}=P\cup\{p_{i}+\Delta p|\Delta p\sim\mathcal{U}(-0.1,0.1),$
+(CCS) $p_{i}\in RS(P,SEV)\}$
 TABLE II: Noise-induced corruptions.
-SEV (severity) refers to the significance of corruption, whose definition varies based on the type of corruption.
-C refers to a constant value, whose concrete value varies in different places.
-RS refers to a random subset function, which samples a subset from the input based on the given $SEV$.
-$\Delta p$ refers to a point in CCS, so $+\Delta p$ means we change the three coordinates in CCS.
-$\Delta r$ refers to the range value, $+\Delta r$ means we only change the range value in SCS.
+SEV (severity) is the significance of corruption, whose definition varies based on the type of corruption.
+C is a constant value, whose concrete value varies in different scenarios.
+RS is a random subset function, which samples a subset from the input based on the given SEVSEV.
+$\Delta p$ is a point in CCS, so $\Delta p$ means we change the three coordinates in CCS.
+$\Delta r$ refers to the range value, $\Delta r$ means we only change the range value in SCS.
 We show the details of noise perturbations in Table II.
@@
 Noises in CCS represent inaccuracies caused by the vibration or rotating of the sensor [38] while SCS can better represent inaccurate measurement of time of flight (ToF) [29].
-In addition to re-positioning existing LiDAR points, we synthesize new points to simulate environment noises, such as dust in the air, through background noise and upsample applied to CCS.
+In addition to re-positioning existing LiDAR points, we synthesize new points to simulate environment noises, such as dust in the air or strong external light, through background noise and upsample applied to CCS.
 Background noise adds points randomly in the entire space, while upsample adds points close to randomly selected points.
@@
 The definitions of these corruptions are listed in Table III.
-III-B Improving Robustness of LiDAR-Based Pose Estimation
-Re-training.
-Learning-based systems are data-driven and could learn to fit themselves into a data distribution.
-We expect these systems to learn how to handle the corrupted data by themselves through re-training, which is a common enhancement and defensive strategy for deep learning-based LiDAR-based algorithms [27, 29, 30].
-Denoising.
-Handcraft algorithms are expected to be tuned and adapted by experts once inefficiencies are identified.
-Therefore such improvements are achieved case by case and lack of generality.
-Hence in this work, for non-learning-based pose estimation systems, we focus on denoising the data, which is a more general way.
-Point cloud denoising removes unwanted noises from LiDAR scans to improve their quality.
-There are three categories of denoising, i.e., filter, optimization, and deep learning-based, each with its advantages and disadvantages [37].
-In this work, we adopted a filter-based method, i.e., the bilateral filter by [40], which is reliable, and significantly faster than the alternatives while preserving the intensity with readily available implementation.
+III-B Improving the Robustness of LiDAR Odometry and Localization
+We propose and experiment with two types of approaches to improve the robustness of such LiDAR systems.
+III-B1 For All LiDAR Odometry and Localization.
+We propose a novel detection-and-filter pipeline to improve the robustness of LiDAR odometry and localization systems.
+This pipeline is algorithm-agnostic and can be widely applied to algorithms of LiDAR pose estimations.
+Our novel pipeline consists of two stages: (1) corruption detection via a DNN-based classifier, and (2) targeted corruption filtering based on the corruption type.
+Stage 1 – Corruption Detection.
+As odometry and localization are real-time applications, we focus on techniques that are inexpensive in terms of time and memory cost.
+Hence, we first convert each 3D LiDAR scan into a 2D range image via spherical projection similar to [7], creating a $64\times 1024$ grid, where each pixel stores the depth and intensity of the point in that cell closest to the agent.
+This representation preserves the spatial structure of the scene while enabling efficient convolutional processing.
+Using these range images, we train a lightweight CNN (Figure 2) to classify data corruptions (e.g., rain, noises) in LiDAR data on our augmented dataset by minimizing categorical cross-entropy loss.
+This loss guides the model to improve classification accuracy across different corruption types through successive training epochs.
+The CNN consists of three convolutional layers with increasing feature sizes (32, 64, and 128), each followed by batch normalization and ReLU activation.
+A final adaptive average pooling layer reduces the output to a 128-dimensional vector to produce class scores, where the class with highest probability is selected as the predicted class for the filtering stage.
+This lightweight architecture is designed for real-time or embedded deployment, balancing classification accuracy with minimal computational overhead.
+Figure 2: Architecture of CNN for corruption classification.
+Stage 2 – Corruption Removal.
+After identifying the corruption type, we apply the corresponding filter to mitigate its effects and restore point-cloud fidelity.
+While our framework supports any paired detector–filter module (e.g., precipitation or density filters), in this work we concentrate on noise-induced corruptions because other point-cloud-based systems are sensitive to noises [29] and denoising methods are mature and effective in removing noises.
+Thus, once the classifier flags a scan as noise-corrupted, we pass it to a bilateral filter [40] for denoising.
+This modular design allows additional filters to be plugged in for other corruption categories.
+III-B2 For Learning-Based LiDAR Odometry and Localization.
+Learning-based approaches inherently adapt to the characteristics of their training data.
+Hence, to defend these systems against the data corruptions, we augment the training set with perturbations shown to degrade performance by our evaluation framework and then fine-tune the pretrained models on this augmented dataset.
+We fine-tune the model using the original training recipe from the released checkpoint until it converges on the augmented dataset.
+This strategy encourages the neural networks to learn invariant features directly from the data distribution, and has been shown to enhance resilience in LiDAR-based systems [25, 27, 28].
 IV Experiment Setup
-IV-A Subject LiDAR-Based Pose Estimation Systems
-TABLE IV: Subject Systems. The errors are reported based on our replication.
-Category Dataset Year Method Error
-Odometry MULLS [24] 2021 Feature-based 0.56%
-KISS-ICP [15] 2023 Direct 0.51%
-Delora [19] 2021 Learning-based 7.78%
-NeRF-LOAM [22] 2023 NeRF-based 1.70%
-Localization LocNDF [21] 2023 NeRF-based 0.059m
-We experiment with four SOTA LiDAR odometry and one LiDAR localization systems, as shown in Table IV.
-We cover one representative work from different types of mainstream point cloud registration methods, i.e., feature-based, direct, and learning-based methods [10].
-We also include two neural distance field-based approaches, which have become popular since the success of the NeRF [20].
+IV-A Subject LiDAR Pose Estimation Systems
+Category Name Year Method
+Odometry MULLS 2021 Feature-based
+KISS-ICP 2023 Direct
+DeLORA 2021 Learning-based
+NeRF-LOAM 2023 Learning-based
+Localization LocNDF 2023 Learning-based
+TABLE IV: Subject LiDAR odometry and localization techniques.
+We evaluate four odometry and one localization systems, as shown in Table IV.
+The subject systems comprise two representative classical methods, MULLS and kiss-icp, and three learning-based approaches, DeLORA, NeRF-LOAM, and LocNDF.
 MULLS [20] is a feature-based LiDAR odometry that uses ground filtering and principal components analysis to extract features from the point cloud.
@@
 We use KITTI Visual Odometry/SLAM Evaluation 2012 (KITTI) [41] to evaluate the four odometry systems, DeLORA, MULLS, KISS-ICP, and NeRF-LOAM.
+The reason is that KITTI is one of the most popular dataset for odometry and all of the selected odometry systems have been evaluated on KITTI.
 KITTI 00 to 10 sequences are released with ground truths.
@@
 (2) sequences 09 and 10 are of moderate sizes which facilitate the large number of perturbations in a reasonable amount of time and computation resources.
-Apollo South-Bay [26] contains completed ADS recordings covering various driving scenarios.
+We use Apollo South-Bay [24] to evaluate LocNDF.
+Apollo South-Bay contains completed ADS recordings covering various driving scenarios.
+LocNDF released the pretrained model and pre-built map on this dataset.
 In this work, we used a subset of ColumbiaPark-3, consisting of 700 testing frames, which is consistent with LocNDF’s pre-trained model [17].
 IV-C Evaluation Metrics
-For the evaluation of the precision of pose estimation, we adopted the relative pose error (RPE) [41].
-RPE is widely used in previous work [10] to measure and compare the performance of pose estimation.
-RPE is calculated as:
-$RPE_{rot}(\mathscr{F})=\frac{1}{|\mathscr{F}|}\sum_{(i,j)\in\mathscr{F}}\angle[(\hat{p}_{j}\ominus\hat{p}_{i})\ominus(p_{j}\ominus p_{i})]$ (2)
-$RPE_{trans}(\mathscr{F})=\frac{1}{|\mathscr{F}|}\sum_{(i,j)\in\mathscr{F}}||(\hat{p}_{j}\ominus\hat{p}_{i})\ominus(p_{j}\ominus p_{i})||_{2}$ (3)
-where $RPE_{trans}$ measures the relative translation error and $RPE_{rot}$ measures the relative rotation error between the estimated pose and the ground truth, $\mathscr{F}$ is a set of frames $(i,j)$, $\hat{p}$ is the estimated pose and $p$ is the ground truth pose, $\ominus$ is the inverse compositional operator [42] and []delimited-[$\angle$[$\cdot$]∠ [ ⋅ ] is the rotation angle.
+For the evaluation of the precision of pose estimation, we adopted the translational relative pose error (RPEtrans[41].
+RPEtrans is widely used in previous work [13] to measure and compare the performance of pose estimation.
+RPEtrans is calculated as:
+[…]
+where $\mathscr{F}$ is a set of frames $(i,j)$, $\hat{p}$ is the estimated pose and $p$ is the ground truth pose, $\ominus$ is the inverse compositional operator [42].
+A pose $p$ consists of a rotation matrix $R$ and a translation vector $t$.
 V Experiment Results
-V-A Evaluating Robustness of LiDAR-Based Pose Estimation
+V-A RobustLOL: Robustness Results of LiDAR Odometry and Localization
 Method.
-We applied our robustness framework to evaluate the subject odometry systems.
+We applied our robustness framework to evaluate the subject odometry and localization systems.
 We first perturb the KITTI and Apollo-SouthBay datasets with the corruptions in Table I.
-Note that NeRF-LOAM is computationally expensive (i.e., requiring at least 24GB GPU memory and days to finish one a single sequence), so we took a statistically-significant sample of frames from KITTI for NeRF-LOAM as a proof of concept.
-Figure 2: Performance of each system on corrupted KITTI LiDAR Odometry Sequence 09 and 10.
-Errors with significantly different magnitudes from the majority are aligned at their magnitude and the actual values are labeled in the figure.
-For example, an error of 12.79% is aligned to “$>$10.0”.
-If multiple errors overlap at the same magnitude, we label the two highest severity levels.
+Note that NeRF-LOAM is computationally expensive (i.e., requiring at least 24 GB GPU memory and several days to finish a single sequence), so we took a statistically-significant sample of frames from KITTI for NeRF-LOAM as a proof of concept.
+Figure 3: Performance of each system on KITTI 09.
+Figure 4: Performance of each system on KITTI 10.
 Results.
-We show the results of RPEtrans in Figures 2.
-Due to the space limitation, we do not present the RPErot in the paper as it is correlated with RPEtrans.
-We share the detailed results of RPErot in our public artifact.
-For the non-learning systems, we find that they have degraded performance on almost all kinds of LiDAR perturbed data compared to the baseline (i.e., un-perturbed data).
-Furthermore, each system is vulnerable to a particular type of perturbation, where their performance drops significantly.
-The error of MULLS increases from below 0.5% to up to 0.8% on most of the corruptions.
-However, the error of MULLS with “local density decrease” (severity 5) increased to more than 12% for KITTI 09 and 18% for KITTI 10.
-Similarly, kiss-icp is especially vulnerable to “background noise”.
-While the error only increased by less than 0.2% in most cases, the error of kiss-icp caused by background noise increases up to 80% (with severity 5 on KITTI 10) from less than 0.6% RPEtrans.
-Moreover, even the lightest background noise (severity 1) causes a significant error increase from less than 0.6% to more than 3% RPEtrans on KITTI 10.
-We observe different results for the learning-based systems, Delora and NeRF-LOAM.
-Delora is vulnerable to most perturbations, especially those related to noise.
-The RPEtrans error increases from less than 10 percent to tens of percent (up to 50%).
-Lastly, NeRF-LOAM is quite different from others.
-We find the results differ very much between the two KITTI sequences, indicating that NeRF-LOAM’s robustness results highly depend on data.
-In KITTI 09, NeRF-LOAM is not sensitive to most perturbations except for one.
-While in KITTI 10, NeRF-LOAM is shown to be highly sensitive in most cases.
-Also, interestingly, in many cases, it shows better robustness with a higher severity than a lower severity.
-For example, for the cutout perturbation on KITTI 10, NeRF-LOAM performs poorly with severity 4, however, the performance is very well with severity 5 which is close to the baseline.
-NeRF-LOAM optimizes a mini-network online for tens of iterations.
-We hypothesize that the poor and unstable results are because NeRF-LOAM is far from being optimized in tens of iterations for these cases.
-For the LiDAR localization system LocNDF, our evaluation shows that it is very robust against all types of corruption.
-The robustness is at the cost of significant upfront investments (in terms of multiple scans of the environment to build the LiDAR map) and its inability to work in unexplored environments.
-Our results show that the SOTA LiDAR localization system, LocNDF, is robust against all point cloud corruptions.
-SOTA LiDAR odometry systems are vulnerable to one or multiple types of point cloud corruptions.
-V-B Improving Robustness of LiDAR-Based Pose Estimation
+Figure 3 and 4 present RPEtrans on KITTI 09 and 10 for all corruptions of severities 1, 3, and 5.
+We omit the results of LocNDF as we observe no performance degradation under any corruption in our experiment.
+We share the complete results in the released artifacts.
+Non-learning systems show overall robustness to most perturbations with each system particularly sensitive to certain corruption types, leading to significant performance drops.
+MULLS is sensitive to local density decrease (local dec) at severity 5 where RPEtrans increases from 0.3% to over 18% and kiss-icp is sensitive to background noise where RPEtrans rises sharply from less than 0.6% on the clean data to over 80%.
+Our investigation shows that local density reduction disrupts key point clusters which MULLS relies on, making its feature extraction unreliable.
+On the other hand, background noise introduces uniformly distributed noisy points which disrupt kiss-icp’s direct point-on-point registration.
+In contrast, learning-based odometry systems, i.e., DeLORA and NeRF-LOAM are more sensitive to most corruptions.
+Notable exceptions include background noise, spherical-coordinate noises (uniform, Gaussian and impulse noises (SCS)), local density increase, and beam deletion, where performance remains relatively stable due to DeLORA’s 2D range image representation.
+During projection, multiple 3D points may map to the same pixel, with only the closest point retained, which significantly limit the effect of these perturbations.
+Thus, background noise appears as mild pixel-level noise, spherical-coordinate noise alters depth without affecting pixel locations, beam deletion has limited impact due to point cloud redundancy, and local density increase adds points that often overlap with existing pixels, preserving image structure.
+In contrast, other corruptions introduce geometric distortions or remove key features, leading to more severe performance drops.
+Finally, we find that NeRF-LOAM shows substantial instability, particularly on KITTI 10.
+In some cases, increasing corruption severity produces lower errors which is counterintuitive.
+To investigate this, we repeated the experiments under each corruption two additional times (three runs in total) with identical inputs.
+The resulting trajectories and final errors vary substantially across runs.
+We attribute this variability to random network initialization and the limited number of online optimization iterations.
+Our systematic evaluation reveals key vulnerabilities in various LiDAR odometry systems.
+V-B Detection-and-Filter Pipeline
+Figure 5: (a) $RPE_{trans}$ with our detection-and-filter pipeline.
+(b) $RPE_{trans}$ of the fine-tuned DeLORA on the augmented dataset.
+Colored bars (light gray, medium gray, dark gray) report the RPE at severity levels 1, 3, and 5 after applying our pipeline or fine-tuning; the hatched white bars with “/” show the corresponding errors without our pipeline or fine-tuning.
 Method.
-We only concern MULLS, kiss-icp, and Delora in this denoising experiment, since LocNDF is robust against all kinds of corruption, while NeRF-LOAM could not produce consistent results between different trials on the same data.
-Firstly, we apply the bilateral filter to denoise the corrupted data and then re-run all the algorithms on the denoised data.
-Furthermore, since Delora uses a learning-based approach, we augment Delora’s training set (i.e., KITTI 00 to 08 sequences) by applying the proposed point cloud corruptions and re-train Delora on the augmented data.
-We focus on denoising the most representative corruption types that affect the algorithms most, i.e., local density decrease (MULLS and Delora), background noise (kiss-icp), uniform noise (Delora), Gaussian noise (Delora), and rain/snow (moderate effects on all systems).
-We denoise these types of corruption at severity 1, 3, and 5.
-We also denoise the clean data to see if the denoising has side effects on clean data.
-Similarly, we fine-tune Delora on the corruptions that affect it most significantly, namely, rain, snow, fog, upsample, uniform noise, Gaussian noise, impulse noise, local density decrease, cutout, and layer deletion.
-We only augmented the training data with these corruptions at severity 5 and fine-tuned Delora for 6 epochs with the default training hyper-parameters.
+We apply the detection-and-filter pipeline on all systems except NeRF-LOAM because it could not produce consistent results between different trials on the exact same data and LocNDF because it is robust against all proposed corruptions.
+We implement the classifier proposed in Section III-B1 using PyTorch 1.13.1+cu117 and train it to recognize eight representative corruptions: rain, snow, fog, uniform noise (CCS), Gaussian noise (CCS), background noise, local density increase, and local density decrease.
+Due to their similar scattering effects, we merge rain and snow into one class; likewise, uniform and Gaussian noise (CCS) are grouped together for their comparable distortion patterns.
+We train the model on KITTI 09 (80% train, 20% validation) for 10 epochs using a batch size of 8, a learning rate of 0.001 with the Adam optimizer.
+We apply it on KITTI 10 and we use the bilateral filter to restore the noise-corrupted PCD.
+All experiments are run on Ubuntu 22.04 LTS with an AMD Ryzen 32-core CPU, 256 GB RAM, and an NVIDIA GeForce RTX 3090 GPU.
 Results.
-Figure 3: (a) Performance on the denoised data. (b) Performance of the retrained model.
-We present the performance of the three systems on the denoised data for KITTI 09 in Figure 3 (a).
-The results on KITTI 10 are similar and available in our replication package.
-For MULLS, local density decrease (local_dec) at severity 5 is the corruption that affects it most significantly.
-We find that the errors even increased by 2.43% after denoising.
-Besides, we find that denoising could reduce the error by 0.26% for Gaussian noise.
-The changes are insignificant for the remaining cases (less than 0.1%).
-For kiss-icp, background noise (bg_noise) is the corruption that affects it most significantly.
-We find that the denoising can reduce the error by 0.2% (severity 1) to 3.22% (severity 5), that is, the negative effect of the corruption is fully eliminated (we observed the same effects on KITTI 10, where the errors at severity 5 are reduced by 80%).
-However, the denoising causes the errors on rain and snow to increase by 0.13% to 0.18%, and on local density decrease at severity 5 to increase by 0.28%.
-The changes are insignificant for the remaining cases.
-For Delora, we find that denoising can reduce its errors in most cases including the baseline, by 0.44% to 16.08%.
-Similar to MULLS and kiss-icp, we find that the errors in local density decrease (local_dec) increased by 2.74%.
-Bilateral filter can effectively mitigate noise-induced corruption in general, and fully eliminate the negative effects of background noise.
-However, it does not help with density-related corruption (local density decrease).
-We present the performance of the Delora after fine-turning for KITTI 09 in Figure 3 (b).
-We find that re-training can bring a much higher improvement than denoising: the errors are reduced by 2.73% to 35.79% for the baseline and all types of corruption.
-However, for KITTI 10, despite the overall effects being similar, we observed a slight increase in the errors for the baseline and a few corruptions.
-It is important to note that our fine-tuning was relatively naive: we only augmented KITTI 00 to 08 with the selected corruptions at severity 5 and used the original learning rate.
+Our classification model achieves an average accuracy of 95.54%, as shown in Table V.
+We notice that clean and fog have a comparably lower accuracy because they are often misclassified as local density increase.
+The reason is that (i) fog has sparse reflections which increase the points density, and (ii) occasional high-reflecting surfaces in clean data also increases points density.
+Weather Noise Density
+Clean Fog R+S U+G Back Inc Dec
+Acc 90.02% 80.83% 98.32% 100% 100% 100% 97.84%
+TABLE V: Classification results for seven class of noises, i.e., Clean, Fog, Rain+Snow (R+S), Uniform+Gaussian Noise (U+G), Background Noise (Back), Local Density Increase (Inc), and Local Density Decrease (Dec)
+We present the RPEtrans of the three systems incorporated with our pipeline on KITTI 10 in Figure 5 (a).
+MULLS is shown to be robust against most noise in the previous section.
+Applying our detection-and-filter pipeline to the noisy data causes negligible performance degradation (less than 0.1%).
+kiss-icp significantly benefits from the detection-and-filter pipeline, with post-denoising errors returning to the clean-data baseline.
+DeLORA also has a significant error reduction after filtering noises in data, however, the errors are still much higher than the baseline.
+Our classifier could classify noise types accurately and our detection-and-filter pipeline could significantly reduce the error on noise-corrupted point cloud data.
+V-C Fine-Tuning
+Method.
+We fine-tune the learning-based system, i.e., DeLORA, on the dataset augmented with the corruptions that affect it most significantly, namely, rain, snow, fog, upsample, uniform noise (CCS), Gaussian noise (CCS), impulse noise (CCS), local density decrease, cutout, and layer deletion.
+We exclude NeRF-LOAM and LocNDF for the same reasons mentioned in Section V-B.
+Results.
+We present the performance of the fine-tuned DeLORA on corrupted KITTI 10 scans (KITTI 09 results are available in the replication package) in Figure 5 (b).
+We find that the errors are reduced by up to 35.79% for all types of corruption after fine-tuning which brings the performance close to the clean data.
+On the baseline, the fine-tuned DeLORA achieves a 2.7% decrease in RPEtrans on KITTI 09 but a 3.3% increase on KITTI 10 compared to the original model.
+It is important to note that our fine-tuning was comparably simple: we only augmented KITTI 00 to 08 with the selected corruptions at severity 5 and used the original learning rate.
 We believe that further improvements could be achieved by a more refined fine-tuning process, such as augmenting the dataset with the corruption at different severities, and adjusting the learning rate and the ratio between clean and corrupted data.
-Re-training with the corruption data is a more effective way to improve the robustness of the learning-based LiDAR pose estimation system (Delora).
+Fine-tuning with the corruption data is an effective way to improve the robustness of the learning-based LiDAR odometry system (DeLORA).
 VI Conclusion
-In this work, we extensively evaluated the robustness of LiDAR-based pose estimation systems, including odometry and localization, against 18 synthetic real-world point cloud corruptions.
-Our findings indicate that all SOTA LiDAR-based odometry systems are particularly vulnerable to certain corruption types.
-We further show that simple bilateral filtering effectively mitigates noise-induced data corruptions, however, density-related corruptions remain challenging.
-Finally, we show that corruption-based data augmentation not only significantly enhances the robustness of learning-based LiDAR odometry systems but also improves their performance on clean data.
+In this work, we propose RobustLOL, a systematic framework for evaluating the robustness of LiDAR-based odometry and localization systems against 18 types of synthetic real-world point cloud corruptions.
+Our evaluation reveals critical vulnerabilities in different types of odometry systems, with performance degradation of up to 80%.
+We further demonstrate that our simple detection-and-filter pipeline can effectively mitigate noise-induced corruptions, restoring performance to near-clean levels.
+Finally, we show that corruption-based data augmentation substantially enhances the robustness of learning-based systems against point cloud corruptions.
+Notably, it also improved clean-data performance on KITTI 09, though a slight decrease was observed on KITTI 10.